### v1

# ABD: Default–Exception Abduction in Finite First-Order Worlds

Abstract
Abduction in knowledge representation is often framed as “explaining away” inconsistencies between a background theory and observations by hypothesizing missing facts or exceptions. Despite decades of KR work on abduction, there are few modern benchmarks that (i) require genuine first-order relational reasoning, (ii) admit unambiguous, solver-checkable verification, and (iii) produce informative error analyses rather than binary right/wrong judgments. We introduce ABD, a family of default–exception abduction tasks over small finite relational worlds. Each instance provides (a) a set of finite structures with observed facts, and (b) a fixed default-like first-order theory that may be violated by those observations. A model must output a first-order abnormality rule $\alpha(x)$ that defines an exception predicate $\mathsf{Ab}(x)\leftrightarrow\alpha(x)$, restoring satisfiability while keeping exceptions sparse.
We formalize three observation regimes with distinct completion semantics.
ABD-Full
assumes closed-world observation.
ABD-Partial
allows unknown atoms under existential completion: $\alpha$ is valid if some completion makes the repaired theory satisfiable, with cost optimized in the best case.
ABD-Skeptical
uses universal completion: $\alpha$ is valid only if the repaired theory is satisfiable under every completion, with cost measured in the worst case.
Because domains are finite, validity and costs are computed via SMT (Z3), enabling exact verification and controlled difficulty. We evaluate ten frontier LLMs on 600 instances spanning all three scenarios and seven default theories. The best models achieve 90-99% training validity, but parsimony gaps of ${\sim}1$$1.5$ extra exceptions per world remain. Holdout evaluation reveals distinct generalization profiles: in ABD-Full and ABD-Partial, the dominant failure is parsimony inflation; in ABD-Skeptical, it is validity brittleness—rules that work on training often break on holdouts—while survivors show smaller gap inflation.
1 Introduction
Default reasoning is a cornerstone of knowledge representation: we often model domains with rules that hold “normally,” while allowing rare exceptions. In classical KR formalisms—default logic (?), circumscription (?), and stable-model semantics (?)—exceptions are not bugs but a feature: they encode domain irregularities without abandoning general rules. From the perspective of abduction (?; ?), exceptions are precisely what we infer when observations conflict with what a default theory predicts.
Modern language and reasoning models can emit syntactically valid logical expressions, including quantified first-order formulas. Can such models propose correct, compact exception rules that repair default theories across multiple relational worlds? Answering this requires an evaluation setting that removes natural-language ambiguity, supports quantifiers, and remains mechanically checkable.
We propose ABD, a benchmark suite for default–exception abduction on finite first-order worlds. Each instance consists of several small relational structures (“worlds”) and a fixed first-order background theory $\Theta$ that uses an abnormality predicate $\mathsf{Ab}$ to block defaults. Rather than providing labeled targets to be matched, we provide only the theory and the observed facts; the model must output a definition of abnormality: a first-order formula $\alpha(x)$, such that $\mathsf{Ab}(x)\leftrightarrow\alpha(x)$ makes all worlds satisfiable. To make the task genuinely abductive, we score parsimony: among valid hypotheses, we prefer those that mark as few abnormal domain elements as possible.
The output $\alpha(x)$ is a unary concept, but its role is not to predict labels. Instead it is a repair policy that negotiates the tension between (i) a theory that expresses default regularities and (ii) concrete relational observations that may violate them. The objective is therefore not to match a provided target extension, but to restore consistency while introducing minimal exceptional structure.
Three Observation Regimes.
We study three scenarios that share the same interface but differ in how missing information is handled:
ABD-Full
(closed-world), ABD-Partial (existential completion: valid if some completion satisfies the repaired theory), and ABD-Skeptical (universal completion: valid only if every completion satisfies the repaired theory). Parsimony mirrors these quantifiers: best-case cost in ABD-Partial, worst-case in ABD-Skeptical. These regimes are formalized in Section 4.
Evaluation Beyond Validity.
Binary validity quickly saturates for stronger models. However, validity is in principle trivial in this task, by labeling many (or all) elements as abnormal. We therefore emphasize cost-based metrics—how close a model comes to a solver-computed lower bound on abnormality count—and size-conditioned variants that group formulas by syntactic complexity (AST size) to discourage degenerate case-splitting.
Main Empirical Findings.
We evaluate ten frontier models across all three regimes. Holdout evaluation reveals two distinct failure modes: parsimony inflation in ABD-Full/Partial (gaps roughly double on fresh worlds) and validity brittleness in ABD-Skeptical (rules that work on training break on holdouts, though survivors show smaller gap inflation). Models differ qualitatively in their performance profiles, pointing to different reasoning strategies. Details are in Section 6.
Contributions.
  • We formalize default–exception abduction on finite first-order worlds with solver-checkable semantics under three observation regimes: ABD-Full (closed-world, fully observed), ABD-Partial (existential completion under missing atoms), and ABD-Skeptical (universal completion with worst-case, robust exception costs).
  • We introduce cost-based scoring for parsimony, including gap-to-lower-bound, gold-relative gap, and size-conditioned analyses by formula complexity.
  • We describe a difficulty-controlled generator that constructs multi-world instances and eliminates shortcut hypotheses via a lightweight, counterexample-guided (CEGIS-like) procedure.
  • We evaluate ten frontier models—Opus-4.6, Grok4, Grok4.1-fast, Gemini-3.1-Pro, Gemini-3-Pro, GPT-5.2, DeepSeek-Reasoner (DSR), Kimi-K2-Thinking, GPT-4o, and Hermes4222Model references: Anthropic Claude Opus 4.6 (?); xAI Grok 4 (?) and Grok 4.1-fast (?); Google Gemini 3 Pro (?) and Gemini 3.1 Pro (?); OpenAI GPT-5.2 (?); DeepSeek-Reasoner (?); Moonshot Kimi K2 Thinking (?); OpenAI GPT-4o (?); Nous Research Hermes 4 (?).—and provide analyses of validity, parsimony gaps, formula complexity, and holdout generalization. Gemini-3.1-Pro was released close to the conference deadline (on February 19, 2026). We chose to include it alongside Gemini-3-Pro: the recent model substantially improves validity while slightly losing in gap metrics.
2 Related Work
Abduction, Defaults, and Non-Monotonic Reasoning.
Abduction is often formalized as selecting hypotheses that, when added to background knowledge, entail observations while satisfying integrity constraints, typically with a preference for parsimonious explanations (?; ?; ?). This connects directly to non-monotonic KR frameworks for defeasible rules and exceptions, including default logic and circumscription (?; ?), and stable-model/answer-set semantics that support explicit minimization principles (including “minimize abnormalities” encodings) (?; ?; ?; ?). Our benchmark adopts a solver-friendly abnormality-predicate encoding with finite grounding, but evaluates more than satisfiability: we quantify parsimony gaps between a model’s proposed exception rule and a solver-computed minimum-exception lower bound, and we test whether the learned exception policy remains valid under held-out worlds.
Abductive Logic Programming and Diagnosis.
Abductive logic programming (ALP) provides a general framework for explanation under integrity constraints and has been developed in both logic-programming and KR traditions (?; ?). Model-based diagnosis similarly treats faults/abnormalities as latent causes chosen to restore consistency, often with minimality criteria and hitting-set structure (?; ?; ?). Our setting differs in interface and emphasis: the hypothesis is an explicit first-order definition of an abnormality predicate (not a set of abduced ground facts), and it must satisfy a multi-world specification, which makes it possible to separate repair-by-case from compact exception structure that generalizes.
Inductive Logic Programming and Relational Learning.
ILP learns relational rules from examples and background knowledge, with classic systems and formalisms for inducing Horn clauses and relational definitions (?; ?; ?; ?; ?; ?). While our hypotheses are symbolic and relational, our tasks are not supervised concept learning: each instance provides observations plus a default theory, and success depends on proposing an exception rule that restores satisfiability while minimizing abnormalities. This shifts the challenge from matching labels to balancing consistency and parsimony under an explicit non-monotonic objective, with exact verification. Synthesizing $\alpha(x)$ from multi-world specifications is also amenable to symbolic ILP and ASP-based learners such as Progol (?) and ILASP (?), which can search a hypothesis space with completeness guarantees relative to a language bias. Our goal is not to position LLMs as replacements for such solvers but to use the task as a diagnostic: because ABD instances have mechanically checkable solutions, the benchmark quantifies how closely neural models approximate the output of an exact combinatorial search, and where they systematically fall short (validity failures, parsimony inflation, holdout brittleness).
Program Synthesis and Solver-Aided Search.
Our tasks can be viewed as constrained synthesis: produce a symbolic artifact that satisfies a solver-checkable specification. This aligns with solver-aided languages and synthesis paradigms such as Sketch, CEGIS, and syntax-guided synthesis (?; ?; ?; ?). A key difference is that our specifications are multi-world and explicitly costed: hypotheses must simultaneously repair multiple structures, and evaluation tracks both validity and the gap between a model’s abnormality cost and a solver-computed lower bound.
Logic and Relational Reasoning Benchmarks.
A broad set of benchmarks probe reasoning with rules, quantifiers, and structured knowledge, often through natural-language interfaces (?; ?; ?). Such settings are valuable but can conflate logical competence with linguistic priors and make intermediate steps hard to verify. Our benchmark instead presents extensional finite structures and uses solver-backed grounding to make correctness and costs exact, enabling diagnostics about exception sparsity, gap above a solver lower bound, and overfitting across held-out worlds.
Mathematical Reasoning and Formal Proof Benchmarks.
Natural-language math benchmarks such as GSM8K and MATH test multi-step reasoning but remain text-mediated (?; ?). Formal proof benchmarks (e.g., CoqGym, HOList, miniF2F, ProofNet) evaluate mechanically verified outputs in proof assistants (?; ?; ?; ?), and large-scale theorem-proving efforts use learning to guide deductive search (?). Our benchmark is complementary: rather than proving theorems, models must synthesize compact exception definitions that reconcile defaults with observations under a minimum-abnormality objective.
Neuro-Symbolic Learning and Large Language Models.
Neuro-symbolic work explores combining learning with symbolic representations and constraint-based verification (?), including approaches that use abduction as a learning signal (?). Recent LLM analyses highlight that models can produce syntactically plausible symbolic outputs while relying on shortcuts or overly long solutions when unconstrained (?; ?). Our evaluation emphasizes solver-verifiable semantics, parsimony-sensitive metrics (cost and gap), and held-out-world testing designed to expose hypotheses that "repair" observed worlds without capturing a stable general exception rule.
Abductive Reasoning in Natural Language.
Abduction has also been studied in NLP as a model of commonsense explanation and narrative understanding, including classic logical accounts of interpretation-as-abduction (?) and benchmark-style tasks for causal/abductive inference (?; ?; ?). These tasks are valuable for modeling everyday explanation but are mediated by language and typically lack a fully explicit world model with exact cost verification. By contrast, our benchmark isolates abductive structure in a formal finite-model setting where validity and costs are mechanically checkable and have a direct interpretation as exception cardinality.
3 Problem Setup
Finite Relational Worlds.
We fix a relational signature $\Sigma=\{P,Q,R,S,=\}$, with unary predicates $P(\cdot),Q(\cdot)$, binary predicates $R(\cdot,\cdot),S(\cdot,\cdot)$, and equality. A world $W$ has a finite domain $D_{W}=\{a_{1},\ldots,a_{n}\}$ ($|D_{W}|\in\{9{-}12\}$; see Table 2) and (possibly partial) interpretations for the predicates in $\Sigma$. In the fully observed setting (ABD-Full), $W$ specifies complete interpretations $P^{W},Q^{W}\subseteq D_{W}$ and $R^{W},S^{W}\subseteq D_{W}\times D_{W}$. In the partial-observation settings (ABD-Partial and ABD-Skeptical), $W$ additionally designates a set of unknown ground atoms $\Omega_{W}$ whose truth values are not observed. In our benchmark, $\Omega_{W}$ ranges over ground atoms of $R$ and $S$; $P$ and $Q$ are always fully observed. In ABD, the world does not include labels for a target concept to be predicted.
Background Theories with Abnormality.
Each instance specifies a fixed first-order theory $\Theta$ over an extended signature $\Sigma_{\mathsf{Ab}}=\Sigma\cup\{\mathsf{Ab}\}$, where $\mathsf{Ab}(\cdot)$ is a unary abnormality predicate. We use $\Theta$ to express default-like constraints that may be blocked by abnormality. A common pattern is an implication schema of the form
$\forall x\;\Big(\mathrm{Ante}(x)\;\wedge\;\neg\mathsf{Ab}(x)\;\rightarrow\;\mathrm{Cons}(x)\Big),$ (1)
where $\mathrm{Ante}$ and $\mathrm{Cons}$ are first-order formulas over $\Sigma$ (often involving existentials over $R,S$). Intuitively, elements satisfying the antecedent are expected to satisfy the consequent unless they are declared abnormal.
Hypotheses as Definitions of Abnormality.
A model outputs a single first-order formula $\alpha(x)$ (one free variable) over a restricted set of allowed predicates (specified per instance). We interpret the prediction as a definition:
$\forall x\;\big(\mathsf{Ab}(x)\leftrightarrow\alpha(x)\big).$ (2)
This definitional view avoids ambiguity about whether $\mathsf{Ab}$ is an independent choice: it is entirely determined by $\alpha$.
Symbolic Output Language.
To support exact evaluation, $\alpha(x)$ is produced in a simple S-expression syntax (atoms, connectives, and first-order quantifiers) that is parsed deterministically. The grammar admits boolean connectives (and, or, not, implies), first-order quantifiers (forall, exists), and atomic predicates from the allowed signature.
Unknown Atoms and Completions.
In partial worlds, each ground atom of $R$ and $S$ is either known true, known false, or unknown. Let $\Omega_{W}$ be the set of unknown atoms in $W$. A completion $\mathcal{C}$ assigns a truth value to every atom in $\Omega_{W}$, yielding a fully specified world $W^{\mathcal{C}}$. We write $\mathrm{KnownFacts}(W)$ for the conjunction of all observed positive facts and observed negative facts in $W$ (treating atoms not listed as true or unknown as observed false). In fully observed worlds, $\mathrm{KnownFacts}(W)$ coincides with $\mathrm{Facts}(W)$. For a completion $\mathcal{C}$, we take $\mathrm{Facts}(W^{\mathcal{C}})$ to be $\mathrm{KnownFacts}(W)$ together with the literals that fix each atom in $\Omega_{W}$ according to $\mathcal{C}$.
Finite Grounding and Solver Checks.
All domains are finite, so satisfiability of $\Theta$ under a predicted $\alpha$ can be checked by grounding quantifiers over $D_{W}$ and translating the result to SMT. In partial worlds, unknown atoms in $\Omega_{W}$ are represented as free Boolean variables in the SMT encoding, so we can quantify over completions by changing which constraints are asserted. Existential-completion validity (ABD-Partial) reduces to a satisfiability check with unknowns left free, whereas skeptical validity (ABD-Skeptical) reduces to the absence of a counterexample completion (checked as a single UNSAT query that asks whether there exists a completion consistent with $\mathrm{KnownFacts}(W)$ under which some grounded axiom fails given $\mathsf{Ab}\leftrightarrow\alpha$). Because costs are cardinalities over a finite domain, we can compute parsimony objectives exactly (or with thresholded SAT checks) using Z3 (?).
We formalize the existential and universal completion semantics as synthesis problems in Section 4.
4 Default–Exception Abduction Tasks
We now define the three ABD task variants as formal synthesis problems. Fix an instance $\mathcal{I}$ with background theory $\Theta$ and training worlds $\mathcal{W}_{\mathrm{tr}}=\{W_{1},\ldots,W_{k}\}$. The learner outputs $\alpha(x)$ and is evaluated by (i) validity and (ii) parsimony.
4.1 ABD-Full: Abduction Under Full Observation
In ABD-Full, each world $W$ provides a complete interpretation of $P,Q,R,S$ under a closed-world assumption for the given domain (any unlisted atom is false). A hypothesis $\alpha$ is valid on a world $W$ if the repaired theory is satisfiable:
$W\models_{\mathrm{ABD}}\alpha\;\Longleftrightarrow\\ \mathrm{SAT}\big(\Theta\wedge\mathrm{Facts}(W)\wedge\forall x(\mathsf{Ab}(x)\leftrightarrow\alpha(x))\big).$ (3)
An $\alpha$ solves an instance iff it is valid on all training worlds.
Cost.
For a valid hypothesis, we define the per-world abnormality count $\mathrm{cost}_{W}(\alpha)=|\{a\in D_{W}:W\models\alpha[a]\}|.$ The instance-level cost is the sum across worlds:
$\mathrm{Cost}(\alpha;\mathcal{W}_{\mathrm{tr}})=\sum_{W\in\mathcal{W}_{\mathrm{tr}}}\mathrm{cost}_{W}(\alpha).$ (4)
Lower-Bound Baseline and Gap.
To measure parsimony, we compare a model’s cost to a solver-computed lower bound obtained by allowing $\mathsf{Ab}$ to be assigned freely (not necessarily definable by a single $\alpha$). Concretely, for each world we compute
$\mathrm{OptCost}(W)=\min_{\mathsf{Ab}\subseteq D_{W}}\;|\mathsf{Ab}|\;\;\text{s.t.}\;\;\mathrm{SAT}\big(\Theta\wedge\mathrm{Facts}(W)\big),$ (5)
and define $\mathrm{OptCost}(\mathcal{W}_{\mathrm{tr}})=\sum_{W\in\mathcal{W}_{\mathrm{tr}}}\mathrm{OptCost}(W)$. We report the gap $\mathrm{Gap}(\alpha)=\mathrm{Cost}(\alpha;\mathcal{W}_{\mathrm{tr}})-\mathrm{OptCost}(\mathcal{W}_{\mathrm{tr}}).$ Because $\mathrm{OptCost}$ relaxes the single-formula constraint and allows $\mathsf{Ab}$ to vary freely by world, it lower-bounds the best definable cost; thus $\mathrm{Gap}(\alpha)$ conservatively overestimates suboptimality. On small domains the bound can often be matched by large extensional case-splits, motivating our AST- and holdout-based diagnostics.
This baseline is nevertheless solver-verifiable, inexpensive compared to full joint optimization, and empirically discriminative across models.
4.2 ABD-Partial: Existential Completion
ABD-Partial introduces missing information about the observed predicates (Section 3). Recall that $\Omega_{W}$ is the set of unknown atoms in $W$ and $W^{\mathcal{C}}$ denotes the completed world under completion $\mathcal{C}$.
Validity Under Existential Completion.
A hypothesis $\alpha$ is valid on $W$ if there exists a completion that makes the repaired theory satisfiable:
$W\models_{\exists\mathrm{comp}}\alpha\;\Longleftrightarrow\\ \exists\mathcal{C}:\mathrm{SAT}\big(\Theta\wedge\mathrm{Facts}(W^{\mathcal{C}})\wedge\forall x(\mathsf{Ab}(x)\leftrightarrow\alpha(x))\big).$ (6)
Cost Optimized Over Completions.
Because $\alpha$ may mention unknown relations, its extension can depend on the completion. We define cost as the minimum abnormality count over completions that satisfy the repaired theory:
$\mathrm{cost}_{W}^{\exists}(\alpha)=\min_{\mathcal{C}}\;|\{a\in D_{W}:W^{\mathcal{C}}\models\alpha[a]\}|\\ \text{s.t.}\;\mathrm{SAT}\!\big(\Theta\wedge\mathrm{Facts}(W^{\mathcal{C}})\wedge\forall x(\mathsf{Ab}(x)\leftrightarrow\alpha(x))\big).$ (7)
Instance cost sums these per-world values, and gaps are defined relative to a completion-aware lower bound that also leaves $\mathsf{Ab}$ unconstrained:
$\mathrm{OptCost}^{\exists}(W)=\min_{\mathcal{C},\,\mathsf{Ab}\subseteq D_{W}}\;|\mathsf{Ab}|\\ \text{s.t.}\;\mathrm{SAT}\big(\Theta\wedge\mathrm{Facts}(W^{\mathcal{C}})\big).$ (8)
As in ABD-Full, the free-$\mathsf{Ab}$ lower bound is conservative (Section 4.1).
4.3 ABD-Skeptical: Universal Completion
ABD-Skeptical retains partial observation but treats unknown atoms conservatively. Intuitively, a hypothesis should repair the default theory not only for some completion of missing facts, but for all completions consistent with what is observed.
Validity Under Universal Completion.
A hypothesis $\alpha$ is skeptically valid on $W$ if the repaired theory is satisfiable for every completion:
$W\models_{\forall\mathrm{comp}}\alpha\;\Longleftrightarrow\\ \forall\mathcal{C}:\;\mathrm{SAT}\big(\Theta\wedge\mathrm{Facts}(W^{\mathcal{C}})\wedge\forall x(\mathsf{Ab}(x)\leftrightarrow\alpha(x))\big).$ (9)
An $\alpha$ solves an instance iff it is valid on all training worlds under this universal-completion semantics.
Worst-Case Cost Over Completions.
For a skeptically valid hypothesis, the per-world cost is the worst-case abnormality count over completions:
$\mathrm{cost}_{W}^{\forall}(\alpha)=\max_{\mathcal{C}}\;|\{a\in D_{W}:W^{\mathcal{C}}\models\alpha[a]\}|.$ (10)
The instance-level cost is the sum across training worlds:
$\mathrm{Cost}^{\forall}(\alpha;\mathcal{W}_{\mathrm{tr}})=\sum_{W\in\mathcal{W}_{\mathrm{tr}}}\mathrm{cost}_{W}^{\forall}(\alpha).$ (11)
Lower-Bound Baseline and Gap.
We lift the free-$\mathsf{Ab}$ lower bound (Section 4.1) to the skeptical setting by taking the worst case over completions:
$\mathrm{OptCost}^{\forall}(W)=\max_{\mathcal{C}}\;\mathrm{OptCost}(W^{\mathcal{C}}),$ (12)
and define $\mathrm{OptCost}^{\forall}(\mathcal{W}_{\mathrm{tr}})=\sum_{W\in\mathcal{W}_{\mathrm{tr}}}\mathrm{OptCost}^{\forall}(W)$. We report the skeptical gap
$\mathrm{Gap}^{\forall}(\alpha)=\mathrm{Cost}^{\forall}(\alpha;\mathcal{W}_{\mathrm{tr}})-\mathrm{OptCost}^{\forall}(\mathcal{W}_{\mathrm{tr}}).$ (13)
This baseline is again conservative in the sense of Section 4.1.
4.4 Gold Rules and Holdout Evaluation
Planted Gold Rules.
Many instances are generated by selecting a “gold” abnormality rule $\alpha^{\star}$ from a controlled template family and constructing worlds for which $\alpha^{\star}$ is valid and nontrivial. We treat $\alpha^{\star}$ as a generator artifact, not as the unique correct answer: models may find different valid rules with lower or higher cost.
Gold-Relative Cost and “Beating Gold.”
To aid analysis and to detect shortcut hypotheses, we report the gap to the planted rule: $\mathrm{Gap}_{\mathrm{gold}}(\alpha)=\mathrm{Cost}(\alpha)-\mathrm{Cost}(\alpha^{\star}).$ Negative values indicate that the model found a rule with fewer exceptions than the planted one. We log such cases and can optionally refine gold rules by searching in a richer hypothesis pool before finalizing a benchmark release.
Size-Conditioned Metrics.
A model can drive cost down by producing large disjunctive case splits tailored to the finite worlds. To separate such behavior from compact reasoning, we report size-conditioned metrics that group formulas by AST size into bins (Table 9) and compare formulas shorter vs. longer than the planted gold rule (Table 10).
Held-Out Worlds.
To probe generalization, we generate k=5k{=}5 holdout worlds per instance from the same world sampler and theory parameters as training, subject to the same per-world acceptance checks (domain size, non-trivial cost floor), but without the adversarial competitor-elimination filter. Holdout is therefore an out-of-filter distribution refresh: matched in sampling distribution but not adversarially hardened. Holdout worlds are pre-generated and cached for reproducibility. We report holdout validity and cost gaps, as well as $\Delta\text{Gap}=\text{H-Gap}-\text{T-Gap}$ (the gap degradation from training to holdout). For ABD-Skeptical, holdouts additionally require that the planted gold rule satisfies universal-completion validity. Table 1 confirms that train and holdout distributions closely match in world counts and per-world gold costs.
Table 1: Train vs Holdout Distribution Match. Mean world counts and per-world gold costs. N counts instances with available holdout worlds (may differ from Table 2 if some instances lack holdouts).
Worlds Gold cost/world
Scenario N Train Holdout Train Holdout
FULL 195 9.0 5.0 1.78 1.88
PARTIAL 241 8.3 5.0 1.68 1.76
SKEPTICAL 162 6.0 5.0 2.05 2.10
5 Dataset Generation
This paper emphasizes controllable difficulty: we want instances where simple exception rules fail, multiple worlds jointly constrain the solution, and models cannot rely on a single brittle shortcut. We therefore generate instances using three ingredients:
(i) Controlled Hypothesis Families.
We draw gold rules from a template library stratified by quantifier depth and syntactic size. Templates range from simple conjunctions (AST $\approx 5$1010) to nested quantified formulas (AST $\approx 15$3030). A diversity gate caps each template at $\leq 15\%$ of instances to avoid dataset monocultures.
(ii) Multi-World Filtering and Parsimony Floors.
Worlds are accepted only if the induced optimization problem is nontrivial (e.g., minimum abnormality counts exceed a floor) and if the planted rule is near-optimal under the chosen baseline. We also filter on aggregate difficulty (e.g., summed opt cost across worlds).
(iii) CEGIS-Like Elimination of Shortcut Competitors.
To prevent trivial competitors (e.g., very small formulas) from matching the gold performance, we maintain a pool of “survivor” formulas and add additional worlds until surviving shortcuts are eliminated. This produces training sets where the version space is smaller and the intended relational structure is more identifiable. The competitor pool includes formulas up to AST size 15, and we iterate until either all competitors are eliminated or a maximum world budget (capped at 12–15 worlds; averages are 6–9 in this release) is reached.
Table 2 summarizes the released benchmark.
Table 2: Abduction Benchmark Summary. (a) Dataset statistics by scenario. (b) Default theories used in the benchmark; each theory defines when $\mathit{Ab}(x)$ blocks a default conclusion.
(a) Dataset Statistics
Scenario N Train Holdout $|D|$ Theories
FULL 195 9.0 5.0 9–11 T1, T2, T3, T4, T5
PARTIAL 243 8.3 5.0 9–11 T1, T2, T3, T4, T5
SKEPTICAL 162 6.0 5.0 10–12 T1, T2, T3, T4, T5, T6, T7
Total 600
(b) Default Theories: $\phi(x)\land\neg\mathit{Ab}(x)\to\psi(x)$
ID Antecedent $\to$ Consequent
T1 $\exists y(R(x,y)\land P(y))$ $\to$ $Q(x)$
T2 $\exists y(R(x,y)\land P(y))$ $\to$ $\exists z(S(x,z)\land Q(z))$
T3 $\exists y(S(x,y)\land P(y))$ $\to$ $\exists z(R(x,z)\land Q(z))$
T4 $\exists y(R(x,y)\land P(y))$ $\to$ $\exists z(S(x,z)\land\forall w(R(z,w)\to P(w)))$
T5 $\exists y(R(x,y)\land P(y))$ $\to$ $\forall z(S(x,z)\to Q(z))$
T6 $P(x)$ $\to$ $\exists y(R(x,y))$
T7 $P(x)$ $\to$ $\forall y(R(x,y)\to Q(y))$
6 Evaluation and Experiments
Table 3: Abduction Training Performance. Gap = normalized extra exceptions above solver lower bound (free-$\mathsf{Ab}$ baseline); GapG = vs gold reference (lower is better). Best values bold.
FULL PARTIAL SKEPTICAL Overall
Model V% Gap GapG V% Gap GapG V% Gap GapG V% Gap GapG
Opus-4.6 100 1.26 0.74 98 1.02 0.55 99 0.86 0.17 99 1.05 0.51
Grok4 88 0.92 0.41 91 1.54 1.06 89 1.08 0.40 89 1.22 0.68
GPT-5.2 85 0.93 0.42 93 1.30 0.82 100 1.49 0.80 92 1.24 0.70
Gemini-3.1 99 1.27 0.76 97 1.41 0.93 98 1.18 0.50 98 1.30 0.76
Gemini-3 77 1.24 0.72 66 1.22 0.77 93 1.25 0.59 77 1.23 0.70
Grok4.1f 94 1.35 0.84 95 1.70 1.23 96 1.46 0.82 95 1.53 1.00
DSR 94 1.45 0.93 96 1.54 1.07 93 1.73 1.05 95 1.56 1.02
Kimi-K2t 57 1.48 0.96 72 1.69 1.22 80 2.46 1.77 70 1.87 1.32
Hermes4 3 2.85 2.30 1 1.78 1.44 4 3.11 2.56 2 2.81 2.29
GPT-4o 9 3.05 2.59 23 3.19 2.68 29 2.87 2.32 20 3.04 2.52
Table 4: Abduction Training Performance by Theory. V% = validity; GapG = gap vs gold (lower is better). T1–T5 aggregate across all three scenarios; T6–T7 are ABD-Skeptical only. Best values per theory bold.
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
Theory V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG V% GapG
T1 100 0.81 92 0.95 97 1.10 97 1.03 97 0.99 97 1.46 95 1.46 74 2.04 3 4.00 16 5.29
T2 99 0.74 92 0.76 93 0.82 99 0.93 80 0.72 98 1.03 95 1.11 73 1.35 2 1.01 16 2.01
T3 99 0.71 90 0.74 88 0.66 98 0.71 71 0.65 95 1.05 97 0.98 69 1.42 1 2.00 28 1.25
T4 99 0.47 86 0.71 89 0.77 97 0.89 61 0.82 91 1.16 93 1.14 58 1.46 0 14 4.66
T5 98 0.16 90 0.60 97 0.61 98 0.56 92 0.54 93 0.81 97 0.87 80 1.00 9 2.75 5 5.17
T6 100 -0.88 86 -0.71 100 -0.88 100 -0.81 93 -0.79 100 -0.61 93 -0.65 64 -0.62 0 93 1.88
T7 96 0.03 92 0.19 100 0.24 100 0.24 96 1.07 96 0.64 88 0.63 84 1.12 4 1.00 60 1.67
[…]
Opus-4.6 98.8% 61.8% 1.05 2.29 +1.01 14.7
Grok4 89.3% 59.4% 1.22 2.32 +0.95 17.7
GPT-5.2 92.3% 61.8% 1.24 2.46 +0.88 22.9
Gemini-3.1 98.0% 69.7% 1.30 2.37 +0.98 12.4
Gemini-3 76.8% 50.0% 1.23 2.13 +0.80 18.9
Grok4.1f 94.8% 81.3% 1.53 2.58 +0.97 11.1
DSR 94.7% 80.2% 1.56 2.66 +0.98 10.8
Kimi-K2t 69.5% 65.0% 1.87 2.82 +0.93 11.5
Hermes4 2.3% 10.5% 2.81 1.33 +1.19 9.0
GPT-4o 19.8% 20.8% 3.04 4.03 +0.99 8.6
6.1 Prompting and Inference Setup
We employ a zero-shot prompting strategy: no benchmark instances or gold formulas appear as in-context examples. Each prompt consists of four components: (1) a system instruction establishing the role (“expert in first-order logic and abductive reasoning”) and requiring JSON output; (2) a scenario-specific task description that defines the observation semantics (closed-world, existential, or universal completion), the S-expression grammar for $\alpha(x)$, the validity/parsimony/simplicity objectives, and illustrative grammar examples; (3) the instance specification, comprising the allowed/forbidden predicates, the theory axioms, and the training worlds with their predicate interpretations (and, for ABD-Partial/Skeptical, unknown atoms); (4) an output format block requesting a single JSON line with the S-expression formula and a natural-language description. The prompt includes structured reasoning guidelines (“identify which objects must be abnormal; look for a pattern; verify against all worlds”) but instructs models to reason internally and emit only the final JSON answer.
All models are queried at temperature $T=0.1$ for near-deterministic outputs. For models offering a thinking mode, we used the provider-default reasoning configuration with 64K token budget; other models were queried in standard generation mode. Each of the 600 benchmark instances is presented independently (single-turn, no multi-round refinement). All results correspond to API snapshots from January–February 2026; model behavior may differ in future versions.
Each model outputs a candidate $\hat{\alpha}(x)$ in S-expression syntax. We compute training validity and exception cost as defined in Section 4, with cost semantics varying by scenario. All gap values reported are normalized by the number of worlds to keep scores comparable across instances.
6.2 Main Results
Table 3 summarizes training performance across models along two axes: (i) validity (does the repaired theory have any model?), and (ii) parsimony (how close is the exception count to a solver lower bound?). The table reports both Gap (vs. the relaxed free-$\mathsf{Ab}$ lower bound; see Section 4.1) and GapG (vs. planted gold rule). Two patterns stand out.
Several frontier models achieve high training validity (above 90%, with the best reaching nearly 99%), while the older baseline GPT-4o remains far behind. This indicates that ABD is not a “syntax-only” benchmark: producing a well-formed formula is easy, but producing a globally valid repair policy across multiple worlds is not.
Parsimony Separates Strong Models.
Among valid predictions, the dominant source of separation is cost. Using normalized gaps (extra exceptions per world), the strongest models cluster around roughly 1.0–1.6 additional abnormalities beyond the lower bound, whereas weaker baselines incur substantially larger gaps. Interpreting the normalized scale: with $\approx 6$1010 training worlds per instance (Table 2), a gap of $\approx 1.1$ corresponds to roughly 771111 extra abnormal elements per instance.
Several models “beat gold” (negative gold-relative gap) on a nontrivial fraction of instances (Table 11). Gold rules are generator anchors, not ground truth; beating gold often correlates with larger AST size. This signal is useful for refining the competitor pool and identifying gold rules that are overly complex, but it does not indicate “winning” the task—parsimony relative to the solver lower bound remains the primary metric.
6.3 Scenario and Theory Breakdown
Table 3 stratifies results by observation regime (FULL, PARTIAL, SKEPTICAL), while Table 4 breaks down performance by background theory. Note that T1–T5 appear in all three scenarios, so their rows aggregate across regimes; T6–T7 are ABD-Skeptical only.
Partial Observation Trades Feasibility for Stability.
ABD-Partial often increases feasibility (higher validity) for some models by allowing existential completions of unknown atoms, but it does not uniformly improve parsimony. In fact, models can remain valid while drifting farther from the lower bound, reflecting that “having more ways to make the world consistent” does not automatically translate into finding parsimonious repairs.
Skeptical Completion Changes the Failure Mode.
ABD-Skeptical requires hypotheses to remain valid under all completions of unknown atoms, not just some favorable one. Training validity under skeptical semantics is not uniformly lower—some models even achieve higher training validity on Skeptical than on Full (Table 3)—but the cost metric differs: models must minimize the worst-case exception count, which penalizes formulas that “get lucky” under particular completions. The main consequence of robust semantics emerges on holdout (Section 6): the primary failure mode shifts from parsimony inflation (in ABD-Full/ABD-Partial) to validity brittleness, where rules that satisfy universal-completion on training worlds break on holdouts.
Model Families Respond Differently to Missingness.
A notable qualitative contrast in Table 3 is that some models benefit from ABD-Partial (higher validity), while others lose validity under partial observation. This suggests that the partial setting genuinely probes reasoning about what could be true rather than merely making the task easier.
Theory Choice Changes Which Shortcuts Exist.
Across theories (Table 4), T5 is consistently easier, while T3/T4 exhibit larger spread across models. A plausible explanation is that theories with stronger relational structure (e.g., nested quantification or additional dependencies in consequents) constrain the exception rule more tightly, reducing the effectiveness of compact heuristics. This theory-conditioned variation is valuable: it yields a natural set of controlled “difficulty knobs” without changing the output language or evaluation protocol. Additionally, T6 exhibits negative GapG values for most frontier models (Table 4), indicating that model predictions frequently beat the planted gold rule; this suggests T6’s gold rules may be overly complex and flags these instances for generator refinement.
6.4 Holdout Generalization
Training worlds in ABD are adversarially filtered to eliminate shortcut competitors; holdout worlds are drawn from the same sampler with the same per-world acceptance checks but without competitor elimination (Section 4.4). Holdout is therefore an out-of-filter distribution refresh, not an i.i.d. split. Because training worlds are harder by construction, holdout validity can exceed training validity for some models; the primary signal is the change in parsimony among rules that survive on both sets.
[…]
Two Distinct Failure Modes.
Table 5 aggregates across all three scenarios and mixes two effects: (i) whether a rule stays satisfiable on holdouts (validity brittleness), and (ii) how parsimonious it remains when satisfiable (parsimony inflation). Table 6 disentangles these by reporting holdout validity conditioned on training validity. These two failure modes manifest differently across scenarios, as we show next.
Table 6: Holdout Validity Conditional on Train Validity (all scenarios). T-Val = train-valid count; T+H = has holdout data; H||T = holdout-valid given train-valid; H%||T = percentage. Best values bold.
Model T-Val T+H H||T H%||T
Opus-4.6 593 591 371 62.8%
Grok4 535 534 350 65.5%
GPT-5.2 554 553 367 66.4%
Gemini-3.1 588 586 417 71.2%
Gemini-3 461 461 290 62.9%
Grok4.1f 542 541 464 85.8%
DSR 568 567 480 84.7%
Kimi-K2t 417 415 379 91.3%
Hermes4 14 14 14 100.0%
GPT-4o 119 119 88 73.9%
Scenario-Level Holdout Analysis.
[…]
Model FULL PARTIAL SKEPTICAL
[…]
Opus-4.6 100% 87% 1.26 2.72 +1.41 98% 50% 1.02 2.12 +0.95 99% 49% 0.86 1.59 +0.25
Grok4 88% 61% 0.92 1.86 +0.96 91% 72% 1.54 2.79 +1.14 89% 38% 1.08 1.85 +0.43
GPT-5.2 85% 55% 0.93 2.73 +1.48 93% 67% 1.30 2.52 +1.03 100% 61% 1.49 2.06 -0.02
Gemini-3.1 99% 87% 1.27 2.60 +1.35 97% 68% 1.41 2.44 +1.00 98% 52% 1.18 1.80 +0.21
Gemini-3 77% 60% 1.24 2.39 +1.16 66% 41% 1.22 2.21 +0.98 93% 52% 1.25 1.66 +0.11
Grok4.1f 94% 86% 1.35 2.67 +1.31 95% 85% 1.70 2.85 +1.14 96% 69% 1.46 1.83 -0.02
DSR 94% 83% 1.45 2.92 +1.44 96% 85% 1.54 2.70 +1.10 93% 70% 1.73 2.23 +0.09
Kimi-K2t 57% 58% 1.48 2.85 +1.46 72% 67% 1.69 2.83 +1.16 80% 70% 2.46 2.78 +0.10
Hermes4 3% 5% 2.85 5.40 +1.42 1% 1% 1.78 3.00 +1.22 4% 31% 3.11 0.47 +0.96
GPT-4o 9% 17% 3.05 5.58 +2.03 23% 16% 3.19 3.31 +0.85 29% 31% 2.87 3.57 +0.83
Theory-Level Holdout Analysis.
[…]
Theory Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 +0.5 (19) +1.5 (11) +0.4 (19) +0.6 (17) +0.5 (16) +0.5 (23) +0.4 (24) +0.2 (26) +0.0 (1) +1.8 (1)
T2 +1.5 (108) +1.1 (102) +1.4 (103) +1.5 (126) +1.2 (80) +1.3 (132) +1.5 (127) +1.5 (105) +2.1 (3) +1.7 (22)
T3 +1.0 (86) +1.1 (81) +1.0 (65) +1.0 (87) +0.7 (54) +1.0 (102) +1.0 (93) +1.0 (74) +2.4 (1) +1.0 (33)
T4 +0.9 (81) +1.0 (81) +0.9 (93) +0.9 (96) +0.9 (60) +1.0 (114) +1.0 (132) +1.0 (84) +0.8 (4)
T5 +0.6 (61) +0.6 (61) +0.3 (67) +0.5 (73) +0.5 (64) +0.5 (67) +0.4 (80) +0.4 (67) +0.9 (8) +1.4 (1)
T6 -0.2 (3) -0.3 (2) +0.5 (4) -0.3 (2) +0.1 (4) +0.4 (7) +0.1 (5) +0.0 (2) +0.2 (12)
T7 +0.3 (13) +0.3 (12) +0.3 (16) +0.3 (16) +0.4 (12) +0.2 (19) +0.4 (19) +0.2 (21) +0.8 (1) +0.4 (15)
6.5 Complexity and Generalization
[…]
FULL
Opus-4.6 98 +1.6 62 +0.6 50 -0.0
Grok4 89 +1.6 69 +0.7 49 +0.7
GPT-5.2 98 +1.7 68 +0.7 19 +0.8
Gemini-3.1 93 +1.6 71 +0.8 100 +0.1
Gemini-3 93 +1.3 60 +1.0 52 +0.6
Grok4.1f 99 +1.5 62 +0.6 86 +0.2
DSR 97 +1.5 51 +0.8 67 +0.2
Kimi-K2t 100 +1.5 86 +0.4 88 +1.8
PARTIAL
Opus-4.6 80 +1.1 28 +0.5 0
Grok4 89 +1.1 19 +1.1 0
GPT-5.2 89 +1.1 39 +0.7 15 +1.0
Gemini-3.1 78 +1.0 38 +0.7
Gemini-3 80 +1.1 32 +0.5 27 +0.6
Grok4.1f 93 +1.2 36 +0.2 0
DSR 89 +1.1 73 +1.2
Kimi-K2t 95 +1.2 31 +0.9
SKEPTICAL
Opus-4.6 73 +0.3 26 +0.1 10 +2.2
Grok4 67 +0.5 26 +0.2 7 -1.0
GPT-5.2 88 -0.1 24 +0.3 12 +1.7
Gemini-3.1 73 +0.2 18 +0.0 33 +0.5
Gemini-3 77 +0.2 23 -0.1 9 -0.9
Grok4.1f 87 -0.0 31 +0.2
DSR 87 +0.1 27 -0.0
Kimi-K2t 97 +0.1 7 +0.5
[…]
Shorter ($\leq$ gold) Longer (>> gold)
Task $\DeltaGap N VΔ\DeltaGap$
FULL 140 92 +1.52 140 49 +0.60
PARTIAL 135 84 +1.17 135 26 +0.48
SKEPTICAL 121 79 -0.00 121 32 +0.33
Overall 396 85 +0.94 396 36 +0.50
Table 11: Gold-Beating Diagnostics (FULL and PARTIAL tasks). BeatsGold% = fraction of valid instances where model cost << gold cost; AvgImpr = mean per-world cost improvement; AvgAST = mean AST size among gold-beaters. Best values bold.
Model BeatsGold% AvgImpr AvgAST N
Opus-4.6 15.9% 0.6 22.5 94
Grok4 11.8% 0.5 29.6 63
GPT-5.2 18.4% 0.5 52.3 102
Gemini-3.1 10.2% 0.6 19.5 60
Gemini-3 11.5% 0.6 24.1 53
Grok4.1f 6.5% 0.7 20.5 35
DSR 4.4% 0.8 16.8 25
Kimi-K2t 3.4% 0.7 16.1 14
Hermes4 0.0% 0
GPT-4o 0.0% 0
[…]
Paired Comparison: Shorter vs. Longer.
[…]
Gold-Beating Instances.
Table 11 shows that some models beat the planted gold rule on a nontrivial fraction of instances, confirming that gold rules are generator anchors (Section 4.4). This diagnostic identifies instances where the planted rule may be suboptimal and feeds back into generator refinement.
6.6 Discussion
Models differ qualitatively in their performance profiles. Some models (e.g., Gemini 3.1 Pro, DeepSeek Reasoner and Grok 4.1 Fast) excel in validity, while others (e.g., Gemini 3 Pro, Grok 4) achieve lower gaps among valid predictions. GPT-5.2 achieves strong overall performance on training worlds but relies on large formulas that often represent case splits. Taken together, the results support three conclusions.
(i) ABD Is Not Solved.
Parsimony remains challenging: the best systems still incur about one extra abnormality per world beyond a solver lower bound, and this gap roughly doubles on holdouts in ABD-Full/Partial.
(ii) Holdout Evaluation Across All Three Regimes Is Essential.
Training metrics alone are misleading because training sets are adversarially filtered. Holdouts reveal that the answer differs by scenario: parsimony inflation dominates in ABD-Full/Partial, while validity brittleness dominates in ABD-Skeptical, suggesting that robust semantics regularizes cost but makes validity harder to preserve.
(iii) Cost- and Size-Aware Reporting Is Necessary.
Binary validity saturates; cost gaps reveal separation; but gap can be driven down by hypothesis complexity. Improved abductive reasoning involves improvement in all three metrics: validity, gap cost, and formula size.
7 Conclusion
We introduced ABD, a solver-checkable benchmark suite for default–exception abduction over finite first-order worlds. In ABD, models must synthesize a first-order abnormality definition $\alpha(x)$ such that $Ab(x)\leftrightarrow\alpha(x)$ repairs a fixed default theory across multiple worlds, while keeping exceptions sparse. We formalized three observation regimes—ABD-Full (closed-world), ABD-Partial (existential completion with best-case cost), and ABD-Skeptical (universal completion with worst-case cost)—and evaluated ten LLMs under exact SMT-based verification.
Across tasks, frontier models often produce valid repairs on training worlds, but ABD remains unsaturated because parsimony does not saturate. Among strong models, the normalized gap to the solver lower bound remains roughly 1.01.01.61.6 extra abnormalities per world; the gap to the planted gold rule is roughly 0.50.51.01.0, confirming systematic over-use of exceptions. The benchmark was calibrated to be moderately challenging for frontier models; harder instances can readily be generated in the future.
Holdout evaluation reveals that generalization is the key remaining challenge, and that its prevailing failure mode depends on the completion semantics. In ABD-Full and ABD-Partial, the primary generalization failure is parsimony inflation: when a rule remains valid on held-out worlds, its cost gap typically increases substantially (often by about one additional exception per world), suggesting that models learn repairs that are tailored to the filtered training worlds and do not transfer as parsimoniously to fresh samples.
In contrast, ABD-Skeptical induces a different generalization profile. For rules that remain skeptically valid, gap inflation is often smaller than in ABD-Full/Partial, consistent with the worst-case objective acting as an implicit regularizer on exception marking. However, the main failure becomes validity brittleness: rules that satisfy the universal-completion criterion on training worlds frequently fail outright on holdouts. Taken together, these outcomes suggest that robust completion semantics shifts difficulty from “how few exceptions can I mark?” to “can I preserve a universally valid repair policy under distribution refresh?”
Finally, the complexity analyses emphasize that performance cannot be summarized by validity and unbudgeted cost. Large formulas can reduce cost on training worlds but become brittle on holdouts, consistent with case-splitting. Moderate-size formulas often provide the best tradeoff. This motivates reporting ABD performance as a multi-objective profile over validity, parsimony, and syntactic budget.
Limitations.
Domains are small to permit exact SMT verification, which also makes extensional case-splitting feasible; we therefore report size-conditioned and held-out-world metrics to diagnose memorization. Our free-$Ab$ baseline relaxes the single-formula constraint, so gap-to-opt is a conservative overestimate. Finally, planted gold rules are generator anchors rather than uniquely optimal targets; models often find valid solutions of lower AST than gold.
Future work.
A natural next step is to use ABD as a training and diagnosis environment: e.g., solver-in-the-loop refinement that supplies counterexample worlds or counterexample completions, and optimization objectives that explicitly trade off parsimony and robustness. Methodologically, it would be valuable to tighten the lower bound (e.g., by approximating the minimum cost under the single-formula constraint), and to scale beyond small worlds via approximate verification or sampling-based evaluation while preserving exact verification. ABD provides a controlled testbed for studying how learned systems acquire (or fail to acquire) portable default–exception first-order rules.
Appendix A Task Prompts
Each benchmark instance is presented to the model as a single user message preceded by a fixed system prompt. The system prompt is shared across all three observation regimes (ABD-Full, ABD-Partial, ABD-Skeptical). The user message consists of:
  1. 1.
    A task description explaining the observation semantics, S-expression grammar, and validity/parsimony objectives;
  2. 2.
    The predicate scope (allowed and forbidden predicates for the hypothesis);
  3. 3.
    The theory (axioms in S-expression form);
  4. 4.
    The training worlds (domain, predicate interpretations, and—for partial/skeptical—unknown atoms);
  5. 5.
    An output format specification requesting JSON with the formula and a natural-language description.
Below we reproduce the system prompt followed by one concrete example prompt for each scenario, generated directly from the benchmark’s prompt-building code.
A.1 System Prompt (All Scenarios)
The following system prompt is prepended to every API call, regardless of scenario.
You are an expert in first-order logic and abductive reasoning.
Your task is to find concise formulas that explain abnormal behavior in logical systems.
Always output valid JSON with the required fields.
A.2 ABD-Full Example Prompt
The following prompt was generated from a benchmark instance with theory TH7 (T5) and 6 training worlds under the closed-world observation regime. All predicate truth values are fully observed; unlisted atoms are false.
# First-Order Logic Abduction (Full Observation)
## Task Overview
You are given a **default reasoning theory** and several finite "worlds" with **full observation**:
- A set of theory axioms involving an abnormality predicate Ab(x)
- Multiple worlds, each with a finite domain and complete predicate interpretations
- Your task: find a formula alpha(x) that defines which objects are "abnormal"
Each world contains:
- A finite domain of objects (named a0, a1, a2, ...)
- Complete interpretations of predicates P(x), Q(x), R(x,y), S(x,y)
**Closed World Assumption**: Only the facts explicitly listed as TRUE are true. Any predicate application not listed should be assumed FALSE.
## Abnormality Semantics
The abnormality predicate Ab(x) is defined by your formula:
- **Ab(x) <-> alpha(x)**: An object a is abnormal if and only if alpha(a) evaluates to TRUE
The theory axioms use Ab(x) to express default rules with exceptions. For example:
- "P(x) AND NOTAb(x) -> Q(x)" means: normally, P-objects have Q, unless they are abnormal
**Your goal**: Find a formula alpha(x) such that:
1. **Validity**: When Ab(x) is defined as alpha(x), all theory axioms are satisfied in every training world
2. **Parsimony**: Among valid formulas, prefer those that minimize the number of abnormal objects
## Scoring / Tie-breaks (important)
- alpha(x) is a SINGLE global formula shared across all worlds.
- Primary objective: **Validity** (must satisfy all axioms in every training world).
- Secondary objective: **Parsimony** -- minimize the total number of abnormal objects across training worlds:
cost = SUM_world |{ a in Domain_world : alpha(a) is TRUE }|
- Tertiary objective: **Simplicity** -- if multiple formulas have the same validity + cost, prefer smaller formulas (lower AST / fewer operators / fewer quantifiers).
## Output Format
You must output your formula in S-expression syntax. The grammar is:
alpha ::= (P x) -- unary predicate applied to variable
| (R x y) -- binary predicate applied to two variables
| (= x y) -- equality of two variables
| (not alpha) -- negation
| (and alpha_1 alpha_2 ...) -- conjunction (2 or more arguments)
| (or alpha_1 alpha_2 ...) -- disjunction (2 or more arguments)
| (forall v alpha) -- universal quantification
| (exists v alpha) -- existential quantification
**Important constraints:**
- Your formula must have exactly one free variable: x
- All other variables must be bound by quantifiers (forall or exists)
- Variable names should be: x (free), y, z, w (bound by quantifiers)
- **Do NOT mention Ab in your formula.**
- **Use ONLY the predicate symbols listed in AllowedAlphaPredicates for this instance.**
- **If your formula uses any predicate in ForbiddenAlphaPredicates, it is invalid.**
- **Do NOT use object names like a0, a1, ... inside alpha(x). Use variables only.**
- Prefer simpler formulas; aim for minimal abnormal set
- **No implication or biconditional**: Express these using other connectives:
- "A implies B" should be written as ‘(or (not A) B)‘
- "A if and only if B" should be written as ‘(and (or (not A) B) (or (not B) A))‘
## Examples of Valid Formulas
**Note**: Examples below are syntactic illustrations; only use predicates that appear in AllowedAlphaPredicates for THIS instance.
- ‘(P x)‘ -- "x is abnormal if x satisfies P"
- ‘(and (P x) (not (Q x)))‘ -- "x is abnormal if x has P but not Q"
- ‘(exists y (and (R x y) (not (P y))))‘ -- "x is abnormal if x has an R-successor without P"
- ‘(not (exists y (R x y)))‘ -- "x is abnormal if x has no R-successors"
- ‘(forall y (or (not (R x y)) (Q y)))‘ -- "x is abnormal if all R-successors of x have Q"
## Evaluation
Your formula will be evaluated on:
1. **Validity**: Does defining Ab(x) <-> alpha(x) satisfy all theory axioms in every world?
2. **Parsimony**: How many objects are marked as abnormal? (fewer is better)
3. **Generalization**: Does the formula also work on held-out test worlds?
**Note**: A trivially true formula like ‘(or (P x) (not (P x)))‘ makes everything abnormal (valid but not parsimonious). A trivially false formula makes nothing abnormal (may violate axioms).
## Problem Instance
**AllowedAlphaPredicates**: ["P", "R"]
**ForbiddenAlphaPredicates**: ["Ab", "Q", "S"]
**Theory ID**: TH7
**Axioms**:
1. ‘(forall x (implies (and (exists y (and (R x y) (P y))) (not (Ab x))) (exists z (and (S x z) (Q z)))))‘
## Training Worlds
### World W0
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9, a10}
**Predicates** (Closed World Assumption: unlisted atoms are false):
- P: {a7, a3, a0, a9}
- Q: {a0, a6, a8}
- R: {(a10, a9), (a8, a2), (a5, a8), (a0, a2), (a7, a9), (a3, a1), (a4, a4), (a0, a8), (a4, a2), (a2, a1), (a0, a10), (a3, a4), (a9, a9), (a1, a1), (a6, a4), (a4, a0), (a4, a1), (a6, a1), (a10, a0), (a9, a7), (a8, a10), (a6, a10), (a5, a1), (a1, a6), (a1, a8), (a6, a5), (a8, a1)}
- S: {(a4, a7), (a10, a7), (a3, a1), (a4, a3), (a0, a0), (a5, a0), (a4, a6), (a8, a3), (a10, a8), (a4, a5), (a2, a3), (a6, a2), (a9, a10), (a8, a9), (a2, a9), (a5, a5), (a7, a9), (a3, a7), (a5, a6)}
### World W1
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9, a10}
**Predicates** (Closed World Assumption: unlisted atoms are false):
- P: {a6, a1}
- Q: {a5, a10, a0, a2}
- R: {(a10, a7), (a4, a1), (a4, a9), (a9, a5), (a10, a8), (a6, a3), (a7, a7), (a10, a9), (a8, a6), (a4, a0), (a6, a9), (a2, a8), (a1, a1), (a2, a0), (a3, a10), (a7, a3), (a0, a5), (a3, a7), (a10, a5), (a0, a0)}
- S: {(a2, a3), (a10, a7), (a7, a1), (a4, a0), (a4, a7), (a5, a9), (a9, a5), (a2, a5), (a7, a9), (a7, a3), (a1, a1), (a6, a4), (a6, a2), (a6, a1), (a2, a1), (a1, a6), (a8, a4), (a7, a0), (a7, a4), (a8, a10)}
### World W2
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Predicates** (Closed World Assumption: unlisted atoms are false):
- P: {a5, a9}
- Q: {a4, a6, a7, a3}
- R: {(a6, a5), (a2, a7), (a4, a8), (a1, a4), (a0, a5), (a5, a5), (a7, a2), (a4, a3), (a3, a4), (a7, a3), (a6, a3), (a5, a7), (a8, a6), (a7, a5), (a9, a1), (a1, a0), (a6, a7), (a8, a1), (a3, a2), (a4, a1), (a0, a0)}
- S: {(a1, a3), (a2, a1), (a7, a9), (a6, a2), (a0, a5), (a2, a4), (a4, a7), (a5, a2), (a6, a7), (a2, a8), (a1, a8), (a0, a8), (a7, a4), (a8, a1), (a7, a7), (a5, a0), (a3, a8)}
### World W3
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Predicates** (Closed World Assumption: unlisted atoms are false):
- P: {a7, a1, a0}
- Q: {a5, a8}
- R: {(a1, a6), (a2, a6), (a8, a3), (a9, a4), (a8, a8), (a9, a6), (a6, a4), (a0, a0), (a3, a2), (a6, a2), (a3, a0), (a5, a1), (a3, a9), (a7, a5), (a7, a8), (a8, a9), (a0, a5), (a2, a3)}
- S: {(a6, a0), (a6, a3), (a4, a8), (a8, a1), (a6, a4), (a9, a8), (a9, a4), (a3, a7), (a0, a3), (a6, a1), (a0, a1), (a1, a5), (a5, a8), (a9, a3), (a3, a8)}
### World W4
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Predicates** (Closed World Assumption: unlisted atoms are false):
- P: {a1, a4, a5, a7}
- Q: {a7, a2, a9}
- R: {(a5, a8), (a1, a6), (a4, a5), (a8, a3), (a9, a9), (a8, a8), (a0, a0), (a5, a6), (a3, a4), (a2, a9), (a6, a9), (a7, a6), (a1, a0)}
- S: {(a7, a6), (a0, a0), (a9, a2), (a1, a8), (a1, a3), (a2, a1), (a2, a3), (a5, a8), (a3, a1), (a1, a4), (a4, a7)}
### World W5
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Predicates** (Closed World Assumption: unlisted atoms are false):
- P: {a8, a3, a7}
- Q: {a3, a6, a4}
- R: {(a8, a8), (a1, a0), (a9, a4), (a5, a8), (a9, a3), (a3, a5), (a0, a4), (a3, a6), (a8, a3), (a5, a1), (a3, a7), (a2, a4), (a3, a2), (a4, a0), (a5, a3), (a1, a3), (a7, a8), (a6, a4), (a9, a0), (a6, a5), (a1, a5)}
- S: {(a8, a8), (a3, a3), (a4, a9), (a7, a8), (a6, a3), (a6, a5), (a6, a6), (a8, a4), (a9, a4), (a9, a2), (a8, a3), (a1, a4), (a0, a5)}
## Your Task
Analyze the theory axioms and training worlds carefully. Find a formula alpha(x) that defines abnormality such that all axioms are satisfied.
**Solve method (reason internally; do NOT output your reasoning):**
1. Understand the theory axioms: What do they require? When might they be violated?
2. For each world, identify which objects would violate an axiom if they were NOT abnormal
3. Look for a pattern: What property characterizes objects that MUST be abnormal?
4. Check for parsimony: Is there a simpler formula that still satisfies all axioms?
5. Formulate your hypothesis as an S-expression formula
6. **Verify**: For each world, check that defining Ab(x) <-> alpha(x) makes all axioms true
**Reminder**: Validity comes first. Only after you have a valid alpha(x), optimize parsimony (fewer abnormal objects), then simplify the formula.
**Example reasoning** (assumes predicates shown are allowed for alpha in this instance):
- Theory: "P(x) AND NOTAb(x) -> Q(x)" (normally, P-objects have Q)
- World: P(a0)=T, Q(a0)=F, P(a1)=T, Q(a1)=T
- Analysis: a0 has P but not Q, so if a0 is not abnormal, the axiom fails
- Result: alpha(x) = (and (P x) (not (Q x))) makes a0 abnormal, satisfying the axiom
**Important**:
- Your formula must make the axioms true in ALL training worlds
- Minimize the number of abnormal objects when possible
- Do NOT mention Ab in your formula
- Use ONLY predicates from AllowedAlphaPredicates; using ForbiddenAlphaPredicates is invalid
- Do NOT use object names like a0, a1, ... inside alpha(x); use variables only
## Output
Your output must be exactly ONE LINE containing ONLY valid JSON with exactly these keys:
- "formula": a single S-expression formula string with one free variable x
- "description": a short plain-English description (one sentence)
Do NOT include code fences, prefixes, or any other text.
Example of correct output:
{"formula":"(and (P x) (not (Q x)))","description":"Objects with P but not Q are abnormal."}
A.3 ABD-Partial Example Prompt
The following prompt was generated from a benchmark instance with theory TH11 (T6) and 6 training worlds under the existential-completion (partial observation) regime. Known true facts and unknown atoms are listed explicitly; unlisted atoms that are not unknown are known false. The model must find a formula valid under at least one completion of unknowns per world.
# First-Order Logic Abduction (Partial Observation)
## Task Overview
You are given a **default reasoning theory** and several finite "worlds" with **partial observation**:
- A set of theory axioms involving an abnormality predicate Ab(x)
- Multiple worlds, each with a finite domain and partially observed predicate interpretations
- Some predicate facts are **unknown** (not observed)
- Your task: find a formula alpha(x) that defines which objects are "abnormal"
Each world contains:
- A finite domain of objects (named a0, a1, a2, ...)
- Known facts: predicates whose truth values have been observed
- Unknown atoms: predicates whose truth values are hidden
**Observation Rules**:
- Atoms listed under "Known Facts" with TRUE values are TRUE
- Atoms listed under "Unknown Atoms" have unknown truth values
- Any atom NOT listed as TRUE or Unknown is FALSE
## Abnormality Semantics
The abnormality predicate Ab(x) is defined by your formula:
- **Ab(x) <-> alpha(x)**: An object a is abnormal if and only if alpha(a) evaluates to TRUE
The theory axioms use Ab(x) to express default rules with exceptions.
**Your goal**: Find a formula alpha(x) such that:
1. **Validity**: For each world, there EXISTS a completion of unknown atoms such that all theory axioms are satisfied when Ab(x) <-> alpha(x)
2. **Parsimony**: Among valid formulas, prefer those that minimize the number of abnormal objects in the best-case completion
## Scoring / tie-breaks (important)
- alpha(x) is a SINGLE global formula shared across all worlds.
- **Completions are per-world**: each world may use a different completion of its unknown atoms.
- Primary objective: **Validity** (for every training world, there exists at least one completion that makes all axioms true).
- Secondary objective: **Parsimony** -- for each world, we consider the best-case completion (most favorable) and count abnormal objects under that completion.
Total cost is the sum of these per-world best-case abnormal counts.
- Tertiary objective: **Simplicity** -- if multiple formulas have the same validity + cost, prefer smaller formulas (lower AST / fewer operators / fewer quantifiers).
## Output Format
You must output your formula in S-expression syntax. The grammar is:
alpha ::= (P x) -- unary predicate applied to variable
| (R x y) -- binary predicate applied to two variables
| (= x y) -- equality of two variables
| (not alpha) -- negation
| (and alpha_1 alpha_2 ...) -- conjunction (2 or more arguments)
| (or alpha_1 alpha_2 ...) -- disjunction (2 or more arguments)
| (forall v alpha) -- universal quantification
| (exists v alpha) -- existential quantification
**Important constraints:**
- Your formula must have exactly one free variable: x
- All other variables must be bound by quantifiers (forall or exists)
- Variable names should be: x (free), y, z, w (bound by quantifiers)
- **Do NOT mention Ab in your formula.**
- **Use ONLY the predicate symbols listed in AllowedAlphaPredicates for this instance.**
- **If your formula uses any predicate in ForbiddenAlphaPredicates, it is invalid.**
- **Do NOT use object names like a0, a1, ... inside alpha(x). Use variables only.**
- Prefer simpler formulas; aim for minimal abnormal set
- **No implication or biconditional**: Express these using other connectives:
- "A implies B" should be written as ‘(or (not A) B)‘
- "A if and only if B" should be written as ‘(and (or (not A) B) (or (not B) A))‘
## Examples of Valid Formulas
**Note**: Examples below are syntactic illustrations; only use predicates that appear in AllowedAlphaPredicates for THIS instance.
- ‘(P x)‘ -- "x is abnormal if x satisfies P"
- ‘(and (P x) (not (Q x)))‘ -- "x is abnormal if x has P but not Q"
- ‘(exists y (and (R x y) (not (P y))))‘ -- "x is abnormal if x has an R-successor without P"
- ‘(not (exists y (R x y)))‘ -- "x is abnormal if x has no R-successors"
- ‘(forall y (or (not (R x y)) (Q y)))‘ -- "x is abnormal if all R-successors of x have Q"
## Evaluation
Your formula will be evaluated on:
1. **Validity**: For each world, does there exist a completion of unknowns such that Ab(x) <-> alpha(x) satisfies all axioms?
2. **Parsimony**: In the best-case completion, how many objects are abnormal? (fewer is better)
3. **Generalization**: Does the formula work on held-out test worlds?
**Key insight**: Unknown atoms can be assigned TRUE or FALSE to help satisfy the axioms. Your formula should work with the known facts while allowing unknowns to be completed favorably.
## Problem Instance
**AllowedAlphaPredicates**: ["P", "Q", "R"]
**ForbiddenAlphaPredicates**: ["Ab", "S"]
**Theory ID**: TH11
**Axioms**:
1. ‘(forall x (implies (and (exists y (and (R x y) (P y))) (not (Ab x))) (exists z (and (S x z) (forall w (implies (R z w) (P w)))))))‘
## Training Worlds
### World W0
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a3, a7, a5, a8}
- Q: {a6, a7, a0}
- R: {(a3, a8), (a0, a8), (a4, a5), (a0, a3), (a4, a4), (a4, a0), (a8, a6), (a4, a8), (a0, a2)}
- S: {(a7, a2), (a4, a0), (a8, a2), (a7, a5), (a3, a6), (a2, a3), (a1, a4), (a2, a4), (a2, a6), (a6, a5)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a0, a0), (a6, a0), (a8, a7), (a8, a1), (a8, a5), (a3, a4), (a4, a7), (a6, a4), (a8, a3)}
- S: {(a2, a8), (a4, a4), (a8, a5), (a5, a5), (a6, a8), (a7, a8)}
### World W1
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a2, a8, a4, a7}
- Q: {a7, a4}
- R: {(a4, a0), (a8, a8), (a3, a4), (a0, a8), (a8, a7), (a5, a2), (a1, a0), (a3, a6), (a3, a1), (a8, a1)}
- S: {(a6, a4), (a5, a6), (a5, a5), (a8, a6), (a8, a4), (a0, a1)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a8, a0), (a5, a8), (a6, a7), (a4, a7), (a3, a0), (a4, a6), (a5, a7), (a3, a7), (a7, a8)}
- S: {(a6, a7), (a7, a0), (a6, a5), (a6, a6), (a1, a8), (a7, a4)}
### World W2
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a0, a7, a4, a8}
- Q: {a6, a1}
- R: {(a1, a8), (a4, a8), (a2, a5), (a6, a4), (a4, a3), (a0, a6), (a3, a6), (a1, a1), (a6, a8)}
- S: {(a3, a3), (a1, a2), (a5, a0), (a6, a4), (a2, a6), (a5, a3), (a5, a2), (a4, a2), (a4, a0), (a4, a6), (a8, a1), (a8, a8), (a3, a0)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a3, a1), (a6, a1), (a5, a4), (a8, a7), (a8, a4), (a1, a7), (a5, a0), (a0, a3), (a5, a5)}
- S: {(a1, a3), (a7, a0), (a0, a5), (a5, a4), (a2, a5), (a7, a2)}
### World W3
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a3, a4, a2, a1, a0}
- Q: {a8, a3, a5, a1}
- R: {(a8, a2), (a3, a7), (a3, a2), (a7, a0), (a0, a3), (a5, a4), (a7, a6), (a1, a0), (a5, a2)}
- S: {(a5, a0), (a7, a8), (a4, a0), (a8, a8), (a1, a4), (a2, a3), (a8, a7), (a0, a0), (a2, a8), (a6, a5), (a2, a5), (a2, a0)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a4, a1), (a1, a1), (a2, a2), (a8, a5), (a7, a2), (a0, a6), (a0, a1), (a0, a4), (a7, a8)}
- S: {(a6, a1), (a0, a5), (a0, a2), (a6, a6), (a5, a1), (a4, a2)}
### World W4
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a4, a0, a8, a3}
- Q: {a0, a8}
- R: {(a5, a0), (a0, a2), (a7, a6), (a6, a4), (a1, a3), (a5, a2), (a2, a0), (a3, a8)}
- S: {(a8, a8), (a1, a3), (a4, a6), (a2, a4), (a2, a5), (a6, a3)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a2, a3), (a8, a1), (a3, a3), (a8, a8), (a4, a6), (a0, a6), (a0, a7), (a1, a0), (a3, a7)}
- S: {(a4, a0), (a7, a7), (a2, a7), (a0, a4), (a4, a7), (a6, a0)}
### World W5
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a2, a1}
- Q: {a2, a7}
- R: {(a0, a4), (a5, a7), (a3, a6), (a0, a6), (a0, a5), (a8, a2), (a2, a5), (a1, a4), (a0, a1), (a2, a8)}
- S: {(a8, a3), (a5, a0), (a5, a5), (a5, a2), (a4, a5), (a5, a8), (a3, a7), (a6, a7)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a6, a7), (a1, a1), (a7, a3), (a5, a0), (a7, a7), (a7, a6), (a5, a1), (a3, a5), (a4, a2)}
- S: {(a6, a3), (a2, a0), (a8, a2), (a5, a4), (a3, a5), (a2, a1)}
## Your Task
Analyze the theory axioms and training worlds carefully, keeping in mind that some facts are unknown.
**Solve method (reason internally; do NOT output your reasoning):**
1. Understand the theory axioms: What do they require? When might they be violated?
2. For each world, note which predicate facts are known vs. unknown
3. Identify which objects would violate an axiom based on KNOWN facts alone
4. Consider how unknown atoms might be completed to satisfy axioms
5. Look for a pattern: Which objects are **forced** to be abnormal by the KNOWN facts (i.e., no completion can make them normal while keeping axioms true)?
6. Formulate your hypothesis as an S-expression formula
7. **Verify**: For each world, check that there exists some completion of unknowns such that Ab(x) <-> alpha(x) satisfies all axioms
**Example reasoning** (assumes predicates shown are allowed for alpha in this instance):
- Theory: "P(x) AND NOTAb(x) -> Q(x)" (normally, P-objects have Q)
- World:
- Known: P(a0)=T
- Known: P(a1)=T
- Unknown: Q(a1)
- Analysis:
- a0 is forced abnormal: since Q(a0)=F is KNOWN because it is not listed to be
true or unknown. No completion can fix Q(a0) to be true; if a0 were normal, the default would fail.
- a1 does NOT need to be abnormal: we can complete Q(a1)=T, making the default true while keeping a1 normal.
- Candidate: alpha(x) = (and (P x) (not (Q x))) marks exactly the "P but not Q" objects abnormal.
Under a favorable completion, Q(a1)=T so a1 is not abnormal, but a0 remains abnormal because Q(a0)=F is fixed.
**Key insight**: Focus on patterns in the KNOWN facts: those listed to be true, or
those not listed to be true or unknown, and therefore known to be false.
Unknown atoms provide flexibility and can be completed to help satisfy axioms and reduce unnecessary abnormality -- but they cannot override known TRUE/FALSE facts.
**Important**:
- Your formula must allow axioms to be satisfied in ALL training worlds (for some completion)
- Minimize the number of abnormal objects when possible
- Do NOT mention Ab in your formula
- Use ONLY predicates from AllowedAlphaPredicates; using ForbiddenAlphaPredicates is invalid
- Do NOT use object names like a0, a1, ... inside alpha(x); use variables only
## Output
Your output must be exactly ONE LINE containing ONLY valid JSON with exactly these keys:
- "formula": a single S-expression formula string with one free variable x
- "description": a short plain-English description (one sentence)
Do NOT include code fences, prefixes, or any other text.
Example of correct output:
{"formula":"(and (P x) (not (Q x)))","description":"Objects with P but not Q are abnormal."}
A.4 ABD-Skeptical Example Prompt
The following prompt was generated from a benchmark instance with theory TH11 (T6) and 5 training worlds under the universal-completion (skeptical) regime. The world format is identical to ABD-Partial but the validity criterion is stricter: the formula must satisfy the axioms under all completions of unknowns, and parsimony is measured by the worst-case abnormal count.
# First-Order Logic Abduction (Skeptical/Universal Completion)
## Task Overview
You are given a **default reasoning theory** and several finite "worlds" with **partial observation**:
- A set of theory axioms involving an abnormality predicate Ab(x)
- Multiple worlds, each with a finite domain and partially observed predicate interpretations
- Some predicate facts are **unknown** (not observed)
- Your task: find a formula alpha(x) that defines which objects are "abnormal"
Each world contains:
- A finite domain of objects (named a0, a1, a2, ...)
- Known facts: predicates whose truth values have been observed
- Unknown atoms: predicates whose truth values are hidden
**Observation Rules**:
- Atoms listed under "Known Facts" with TRUE values are TRUE
- Atoms listed under "Unknown Atoms" have unknown truth values
- Any atom NOT listed as TRUE or Unknown is FALSE
## Abnormality Semantics
The abnormality predicate Ab(x) is defined by your formula:
- **Ab(x) <-> alpha(x)**: An object a is abnormal if and only if alpha(a) evaluates to TRUE
The theory axioms use Ab(x) to express default rules with exceptions.
## SKEPTICAL Validity (IMPORTANT DIFFERENCE)
Unlike partial observation with existential completion, this task uses **skeptical (universal) completion**:
**Your goal**: Find a formula alpha(x) such that:
1. **Skeptical Validity**: For each world, **FOR ALL completions** of unknown atoms, the theory axioms are satisfied when Ab(x) <-> alpha(x)
2. **Parsimony**: Among valid formulas, prefer those that minimize the **WORST-CASE** number of abnormal objects
Your formula must be **robust** -- it must ensure the axioms hold regardless of how the unknown atoms are resolved.
## Scoring / tie-breaks (important)
- alpha(x) is a SINGLE global formula shared across all worlds.
- **Completions are per-world**: each world considers ALL possible completions of its unknown atoms.
- Primary objective: **Skeptical Validity** (for every training world, ALL completions must satisfy all axioms).
- Secondary objective: **Parsimony** -- for each world, we consider the worst-case completion (least favorable) and count abnormal objects under that completion.
Total cost is the sum of these per-world worst-case abnormal counts.
- Tertiary objective: **Simplicity** -- if multiple formulas have the same validity + cost, prefer smaller formulas (lower AST / fewer operators / fewer quantifiers).
## Output Format
You must output your formula in S-expression syntax. The grammar is:
alpha ::= (P x) -- unary predicate applied to variable
| (R x y) -- binary predicate applied to two variables
| (= x y) -- equality of two variables
| (not alpha) -- negation
| (and alpha_1 alpha_2 ...) -- conjunction (2 or more arguments)
| (or alpha_1 alpha_2 ...) -- disjunction (2 or more arguments)
| (forall v alpha) -- universal quantification
| (exists v alpha) -- existential quantification
**Important constraints:**
- Your formula must have exactly one free variable: x
- All other variables must be bound by quantifiers (forall or exists)
- Variable names should be: x (free), y, z, w (bound by quantifiers)
- **Do NOT mention Ab in your formula.**
- **Use ONLY the predicate symbols listed in AllowedAlphaPredicates for this instance.**
- **If your formula uses any predicate in ForbiddenAlphaPredicates, it is invalid.**
- **Do NOT use object names like a0, a1, ... inside alpha(x). Use variables only.**
- Prefer simpler formulas; aim for minimal worst-case abnormal set
- **No implication or biconditional**: Express these using other connectives:
- "A implies B" should be written as ‘(or (not A) B)‘
- "A if and only if B" should be written as ‘(and (or (not A) B) (or (not B) A))‘
## Examples of Valid Formulas
**Note**: Examples below are syntactic illustrations; only use predicates that appear in AllowedAlphaPredicates for THIS instance.
- ‘(P x)‘ -- "x is abnormal if x satisfies P"
- ‘(and (P x) (not (Q x)))‘ -- "x is abnormal if x has P but not Q"
- ‘(exists y (and (R x y) (not (P y))))‘ -- "x is abnormal if x has an R-successor without P"
- ‘(not (exists y (R x y)))‘ -- "x is abnormal if x has no R-successors"
- ‘(forall y (or (not (R x y)) (Q y)))‘ -- "x is abnormal if all R-successors of x have Q"
- ‘(exists y (exists z (and (R x y) (R x z) (not (= y z)))))’ - "x has at least two distinct R-successors"
## Evaluation
Your formula will be evaluated on:
1. **Skeptical Validity**: For each world, do ALL completions of unknowns lead to Ab(x) <-> alpha(x) satisfying all axioms?
2. **Parsimony**: In the worst-case completion, how many objects are abnormal? (fewer is better)
3. **Generalization**: Does the formula work on held-out test worlds?
**Key insight**: Your formula must be robust enough that no matter how the unknown atoms are resolved, the axioms still hold. This is more demanding than existential completion -- you cannot rely on a "helpful" assignment of unknowns.
## Problem Instance
**AllowedAlphaPredicates**: ["P", "Q", "R"]
**ForbiddenAlphaPredicates**: ["Ab", "S"]
**Theory ID**: TH11
**Description**: existsy(R(x,y)ANDP(y)) AND NOTAb(x) -> existsz(S(x,z)ANDforallw(R(z,w)->P(w))): Nested universal in consequent
**Axioms**:
1. ‘(forall x (implies (and (exists y (and (R x y) (P y))) (not (Ab x))) (exists z (and (S x z) (forall w (implies (R z w) (P w)))))))‘
## Training Worlds
### World W0
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a3, a8}
- Q: {a5, a6, a7, a8, a9}
- R: {(a0, a0), (a0, a5), (a0, a6), (a1, a4), (a1, a7), (a2, a3), (a2, a4), (a2, a5), (a2, a8), (a3, a1), (a3, a3), (a3, a4), (a5, a4), (a5, a6), (a5, a8), (a6, a4), (a7, a3), (a7, a4), (a7, a5), (a7, a6), (a7, a9), (a8, a5), (a8, a6), (a8, a8), (a8, a9), (a9, a1), (a9, a4), (a9, a9)}
- S: {(a0, a4), (a1, a1), (a1, a2), (a1, a5), (a3, a8), (a4, a6), (a4, a7), (a5, a1), (a5, a4), (a5, a7), (a5, a8), (a6, a4), (a6, a5), (a7, a3), (a7, a4), (a7, a6), (a8, a0), (a8, a1), (a8, a8), (a9, a2), (a9, a3), (a9, a4), (a9, a6), (a9, a7), (a9, a8)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a8, a3), (a2, a1), (a6, a3), (a3, a8), (a6, a9)}
- S: {(a6, a8), (a7, a9), (a6, a2), (a6, a9), (a5, a6)}
### World W1
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a4, a7}
- Q: {a3, a4}
- R: {(a0, a2), (a0, a6), (a1, a6), (a2, a5), (a2, a6), (a3, a1), (a3, a6), (a3, a8), (a3, a9), (a4, a0), (a4, a3), (a4, a4), (a4, a5), (a4, a6), (a4, a8), (a4, a9), (a5, a3), (a5, a5), (a5, a6), (a7, a1), (a7, a5), (a7, a6), (a8, a3), (a8, a5), (a8, a6), (a9, a3), (a9, a4), (a9, a5), (a9, a6), (a9, a7)}
- S: {(a0, a0), (a0, a1), (a0, a2), (a0, a5), (a0, a6), (a0, a7), (a0, a8), (a0, a9), (a1, a1), (a1, a2), (a1, a3), (a1, a4), (a1, a5), (a1, a6), (a1, a7), (a1, a8), (a1, a9), (a2, a0), (a2, a1), (a2, a2), (a2, a3), (a2, a6), (a4, a9), (a9, a6)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a4, a1), (a1, a9), (a8, a8)}
- S: {(a7, a0), (a5, a1), (a9, a4), (a4, a3), (a5, a7), (a7, a1), (a2, a4)}
### World W2
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a4, a7}
- Q: {a3, a4}
- R: {(a0, a2), (a0, a6), (a1, a6), (a2, a5), (a2, a6), (a3, a1), (a3, a6), (a3, a8), (a3, a9), (a4, a0), (a4, a3), (a4, a5), (a4, a6), (a4, a8), (a4, a9), (a5, a3), (a5, a5), (a5, a6), (a7, a1), (a7, a5), (a7, a6), (a8, a3), (a8, a5), (a8, a6), (a9, a3), (a9, a4), (a9, a5), (a9, a6), (a9, a7)}
- S: {(a0, a0), (a0, a1), (a0, a2), (a0, a5), (a0, a6), (a0, a7), (a0, a8), (a0, a9), (a1, a1), (a1, a2), (a1, a3), (a1, a5), (a1, a6), (a1, a7), (a1, a8), (a1, a9), (a2, a0), (a2, a1), (a2, a2), (a2, a3), (a2, a4), (a2, a6), (a4, a9), (a9, a6)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a4, a4), (a8, a9), (a0, a0), (a8, a2), (a3, a2)}
- S: {(a1, a4), (a5, a8), (a6, a7), (a3, a9), (a7, a7)}
### World W3
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a4, a7}
- Q: {a3, a4}
- R: {(a0, a2), (a0, a6), (a1, a6), (a2, a5), (a2, a6), (a3, a1), (a3, a6), (a3, a8), (a3, a9), (a4, a0), (a4, a3), (a4, a4), (a4, a5), (a4, a6), (a4, a8), (a4, a9), (a5, a3), (a5, a5), (a5, a6), (a7, a1), (a7, a5), (a8, a3), (a8, a5), (a8, a6), (a9, a3), (a9, a4), (a9, a5), (a9, a6), (a9, a7)}
- S: {(a0, a0), (a0, a1), (a0, a2), (a0, a5), (a0, a6), (a0, a7), (a0, a8), (a0, a9), (a1, a1), (a1, a3), (a1, a4), (a1, a5), (a1, a6), (a1, a7), (a1, a8), (a1, a9), (a2, a0), (a2, a1), (a2, a2), (a2, a3), (a2, a4), (a2, a6), (a4, a9), (a9, a6)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a5, a2), (a7, a6)}
- S: {(a1, a2), (a8, a0), (a6, a9), (a8, a4), (a3, a2), (a5, a8), (a5, a2), (a5, a4)}
### World W4
Domain: {a0, a1, a2, a3, a4, a5, a6, a7, a8, a9}
**Known Facts** (unlisted atoms that are not Unknown are known FALSE):
- P: {a4, a7}
- Q: {a3, a4}
- R: {(a0, a2), (a0, a6), (a1, a6), (a2, a5), (a2, a6), (a3, a1), (a3, a6), (a3, a8), (a3, a9), (a4, a0), (a4, a3), (a4, a4), (a4, a5), (a4, a6), (a4, a8), (a4, a9), (a5, a3), (a5, a5), (a5, a6), (a7, a1), (a7, a5), (a7, a6), (a8, a3), (a8, a5), (a8, a6), (a9, a3), (a9, a4), (a9, a5), (a9, a6), (a9, a7)}
- S: {(a0, a0), (a0, a1), (a0, a2), (a0, a6), (a0, a7), (a0, a8), (a0, a9), (a1, a2), (a1, a3), (a1, a4), (a1, a5), (a1, a6), (a1, a7), (a1, a8), (a1, a9), (a2, a0), (a2, a1), (a2, a2), (a2, a3), (a2, a4), (a2, a6), (a4, a9), (a9, a6)}
**Unknown Atoms** (truth value not observed, can be completed either way):
- R: {(a2, a1), (a7, a0), (a4, a2)}
- S: {(a3, a1), (a3, a9), (a7, a4), (a9, a5), (a0, a5), (a1, a1), (a3, a6)}
## Your Task
Analyze the theory axioms and training worlds carefully. Remember: your formula must work for ALL possible completions of unknown atoms, not just some favorable completion.
**Solve method (reason internally; do NOT output your reasoning):**
1. Understand the theory axioms: What do they require? When might they be violated?
2. For each world, note which predicate facts are known vs. unknown
3. Identify which objects would violate an axiom in ANY completion
4. Consider the WORST-CASE completion: which objects might be forced abnormal when unknowns are resolved unfavorably?
5. Look for a pattern: Which objects must be abnormal to ensure axioms hold NO MATTER HOW unknowns are completed?
6. Formulate your hypothesis as an S-expression formula
7. **Verify**: For each world, check that for ALL completions of unknowns, Ab(x) <-> alpha(x) satisfies all axioms
**Example reasoning** (assumes predicates shown are allowed for alpha in this instance):
- Theory: "P(x) AND NOTAb(x) -> Q(x)" (normally, P-objects have Q)
- World:
- Known: P(a0)=T
- Known: P(a1)=T
- Unknown: Q(a1)
- Analysis for SKEPTICAL semantics:
- a0 is forced abnormal: Q(a0)=F is KNOWN, so if a0 were normal, the default would fail.
- a1 MUST ALSO be abnormal under skeptical semantics: Q(a1) is unknown, so in the WORST-CASE completion where Q(a1)=F, the default would fail if a1 were normal.
- Unlike existential completion where we could rely on Q(a1)=T, skeptical requires robustness against Q(a1)=F.
- Candidate: alpha(x) = (P x) marks ALL P-objects abnormal, ensuring robustness.
**Key insight**: Focus on what might go WRONG if unknowns are completed adversarially:
- An object with unknown predicates in the antecedent/consequent of a default might be forced abnormal to handle the worst case.
- You cannot assume a "helpful" completion of unknowns -- you must be prepared for the least favorable assignment.
**Important**:
- Your formula must make axioms satisfied in ALL training worlds for ALL completions
- Minimize the worst-case number of abnormal objects when possible
- Do NOT mention Ab in your formula
- Use ONLY predicates from AllowedAlphaPredicates; using ForbiddenAlphaPredicates is invalid
- Do NOT use object names like a0, a1, ... inside alpha(x); use variables only
## Output
Your output must be exactly ONE LINE containing ONLY valid JSON with exactly these keys:
- "formula": a single S-expression formula string with one free variable x
- "description": a short plain-English description (one sentence)
Do NOT include code fences, prefixes, or any other text.
Example of correct output:
{"formula":"(P x)","description":"All P-objects are abnormal to ensure robustness against any completion."}
Appendix B Output Language Specification
B.1 S-Expression Grammar
Models must output hypotheses in the following S-expression syntax (EBNF). The grammar is shared across all three observation regimes.
alpha ::= atom
| "(not" alpha ")"
| "(and" alpha alpha+ ")"
| "(or" alpha alpha+ ")"
| "(forall" var alpha ")"
| "(exists" var alpha ")"
atom ::= "(" pred var ")" (* unary *)
| "(" pred var var ")" (* binary *)
| "(=" var var ")" (* equality *)
pred ::= "P" | "Q" | "R" | "S"
var ::= "x" | "y" | "z" | "w"
Implication and biconditional are not primitive; models are instructed to encode them:
  • $A\to B$ as (or (not A) B)
  • $A\leftrightarrow B$ as (and (or (not A) B) (or (not B) A))
Variable discipline. The hypothesis $\alpha(x)$ must contain exactly one free variable $x$; all other variables must be bound by quantifiers, and no object constants (a0, a1, …) may appear in $\alpha$.
B.2 Predicate Scoping Rules
Each benchmark instance specifies two predicate sets:
AllowedAlphaPredicates:
The predicates that $\alpha(x)$ may reference. These always form a subset of $\{P,Q,R,S\}$ and vary by theory (see Appendix C).
ForbiddenAlphaPredicates:
Predicates that must not appear in $\alpha(x)$. This set always includes $\mathsf{Ab}$ (since $\alpha$ defines $\mathsf{Ab}$) and also includes the repaired predicates—those appearing in the consequent of the default rule—to prevent trivial shortcut solutions such as $\alpha(x)=\lnot Q(x)$.
Additional constraints. The hypothesis $\alpha(x)$ must have exactly one free variable $x$. All other variables must be bound by quantifiers. Object names (a0, a1, …) may not appear inside $\alpha$; only logical variables are permitted.
B.3 Complexity Measures
We use two standard measures of formula complexity throughout the paper.
AST Size ($|\alpha|$).
The number of nodes in the abstract syntax tree of $\alpha$:
  • Each predicate application counts as $1+$ the number of its arguments (each argument variable counts as 1).
  • Equality counts as $1+2=3$.
  • Negation: $1+|\varphi|$.
  • Binary connectives ($\land$, $\lor$): $1+|\varphi_{1}|+|\varphi_{2}|$.
  • Quantifiers ($\forall v$, $\exists v$): $1\text{ (quantifier)}+1\text{ (bound variable)}+|\varphi|$.
For example, (exists y (and (R x y) (P y))) has AST size =2+(1+(3+2))=8=2+(1+(3+2))=8: exists(1) + y(1) + and(1) + R(1) + x(1) + y(1) + P(1) + y(1).
Quantifier Depth ($\mathrm{QD}(\alpha)$).
The maximum nesting depth of quantifiers, ignoring connectives:
  • Atomic formulas and equality: 0.
  • Negation, conjunction, disjunction: inherited from children ($\max$ for binary connectives).
  • $\forall v.\,\varphi$ or $\exists v.\,\varphi$: $1+\mathrm{QD}(\varphi)$.
For example, (P x) has $\mathrm{QD}=0$; (exists y (R x y)) has $\mathrm{QD}=1$; (forall y (exists z (R y z))) has $\mathrm{QD}=2$.
Appendix C Theory Library
The benchmark uses seven default-exception theories of the form $\forall x\bigl(\phi(x)\land\lnot\mathsf{Ab}(x)\to\psi(x)\bigr)$, where $\phi$ is the antecedent (trigger condition) and $\psi$ is the consequent (default conclusion). Theories T1–T5 are used in all three scenarios; T6–T7 are used only in ABD-Skeptical.
C.1 Theory Formulas
Table C.1 lists each theory with its full S-expression axiom.
Table C.1: Theory library. Each theory is a single default rule $\phi(x)\land\lnot\mathsf{Ab}(x)\to\psi(x)$.
ID Antecedent $\phi(x)$ Consequent $\psi(x)$
T1 $\exists y\bigl(R(x,y)\land P(y)\bigr)$ $Q(x)$
T2 $\exists y\bigl(R(x,y)\land P(y)\bigr)$ $\exists z\bigl(S(x,z)\land Q(z)\bigr)$
T3 $\exists y\bigl(S(x,y)\land P(y)\bigr)$ $\exists z\bigl(R(x,z)\land Q(z)\bigr)$
T4 $\exists y\bigl(R(x,y)\land P(y)\bigr)$ $\exists z\bigl(S(x,z)\land\forall w(R(z,w)\to P(w))\bigr)$
T5 $\exists y\bigl(R(x,y)\land P(y)\bigr)$ $\forall z\bigl(S(x,z)\to Q(z)\bigr)$
T6 $P(x)$ $\exists y\bigl(R(x,y)\bigr)$
T7 $P(x)$ $\forall y\bigl(R(x,y)\to Q(y)\bigr)$
C.2 Allowed and Forbidden Predicates
Table C.2 shows, for each theory, which predicates may appear in the hypothesis $\alpha(x)$ and which are forbidden. Forbidden predicates include $\mathsf{Ab}$ (always) plus the repaired predicates—those appearing in the consequent—to block trivial shortcut solutions.
Table C.2: Predicate scoping per theory. Repaired = consequent predicates excluded from $\alpha$ to prevent shortcuts.
ID Allowed in $\alpha$ Forbidden in $\alpha$ Repaired
T1 $P,R,S$ $\mathsf{Ab},Q$ $Q$
T2 $P,R$ $\mathsf{Ab},S,Q$ $S,Q$
T3 $P,S$ $\mathsf{Ab},R,Q$ $R,Q$
T4 $P,Q,R$ $\mathsf{Ab},S$ $S$
T5 $P,R,S$ $\mathsf{Ab},Q$ $Q$
T6 $P,Q,S$ $\mathsf{Ab},R$ $R$
T7 $P,R,S$ $\mathsf{Ab},Q$ $Q$
C.3 Theory Commentary
T1
(relational antecedent, unary consequent). The simplest relational theory; checking the default reduces to verifying $Q(x)$ for each domain element triggered by an $R$-neighbor with $P$. Models frequently achieve near-optimal cost.
T2
[…]
T3
(swapped relations: $S$-antecedent, $R$-consequent). Mirrors T2 with $S$ and $R$ exchanged. Since $R$ is repaired (forbidden in $\alpha$), the hypothesis space is restricted to $\{P,S\}$, which limits the available structural patterns and often forces models toward simple unary heuristics.
T4
(nested universal in consequent). The consequent demands an $S$-witness $z$ all of whose $R$-successors satisfy PP—a second-order-flavored constraint. This makes validity checking expensive and produces the highest syntactic complexity in gold rules.
T5
(universal consequent over $S$). Requires that all $S$-successors of the non-abnormal object have $Q$. A single counterexample (an $S$-successor without $Q$) is enough to trigger abnormality, giving models a relatively readable signal.
T6
(unary antecedent, existential consequent; Skeptical-only). The simplest consequent structure: $P(x)$ just requires some $R$-successor. Model predictions frequently beat the planted gold rule (negative GapG), suggesting that the gold formulas are overly complex relative to the constraint.
T7
[…]
Appendix D Dataset Generation Details
This section describes the generation pipeline as pseudocode and parameter tables. All parameters correspond to the released benchmark (abd_b1_v1).
D.1 World Sampler
Each world is a finite first-order structure over four predicates ($P$, $Q$ unary; $R$, $S$ binary) plus the derived $\mathsf{Ab}$.
Domain size.
Domain sizes vary by scenario. In ABD-Full and ABD-Partial, each world’s domain is drawn uniformly from $\{9,10,11\}$; different worlds within the same instance may have different sizes (and in practice nearly all instances span the full 9–11 range). In ABD-Skeptical, domains range over $\{10,11,12\}$; all worlds within a single instance share the same domain size. Objects are named $a_{0},a_{1},\ldots$
Predicate densities.
For each world the sampler draws a density $\rho$ per predicate uniformly from the ranges in Table D.1, then marks $\max(1,\lfloor n^{k}\cdot\rho\rfloor)$ ground atoms true (chosen uniformly at random), where $n$ is the domain size and $k$ is the arity.
Table D.1: Predicate density ranges by scenario. For each world the sampler draws $\rho$ uniformly from the listed interval; $n=|D|$. ABD-Full and ABD-Partial share the same ranges; ABD-Skeptical uses denser binary predicates and a narrower $P$ range.
Predicate Arity Full / Partial Skeptical True atoms
$P$ 1 [0.20, 0.60][0.20,\;0.60] [0.40, 0.60][0.40,\;0.60] $\lfloor n\cdot\rho_{P}\rfloor$
$Q$ 1 [0.20, 0.60][0.20,\;0.60] [0.20, 0.50][0.20,\;0.50] $\lfloor n\cdot\rho_{Q}\rfloor$
$R$ 2 [0.12, 0.25][0.12,\;0.25] [0.15, 0.30][0.15,\;0.30] $\lfloor n^{2}\cdot\rho_{R}\rfloor$
$S$ 2 [0.08, 0.18][0.08,\;0.18] [0.10, 0.25][0.10,\;0.25] $\lfloor n^{2}\cdot\rho_{S}\rfloor$
Unknown atoms (ABD-Partial / Skeptical).
After sampling a complete world, a fraction of $R$ and $S$ atoms are masked as unknown. Table D.2 lists the per-theory masking rates used for each scenario. Unary predicates $P$, $Q$ are never masked.
Table D.2: Unknown-atom rates (fraction of binary atoms masked). ABD-Partial uses the default column; ABD-Skeptical uses theory-specific rates to keep universal validity tractable.
ABD-Partial ABD-Skeptical
Theory $R$ $S$ $R$ $S$
T1 (TH2) 0.20 0.10 0.05 0.08
T2 (TH7) 0.20 0.10 0.05 0.05
T3 (TH10) 0.20 0.10 0.05 0.05
T4 (TH11) 0.20 0.10 0.05 0.05
T5 (TH12) 0.20 0.10 0.05 0.05
T6 (TH3) 0.04 0.08
T7 (TH5) 0.05 0.08
D.2 Filtering Criteria
Every candidate instance must pass the following filters before inclusion.
Validity floor.
The gold hypothesis $\alpha^{*}$ must be valid on all training worlds (all scenarios) and achieve $\mathrm{opt\_cost}(\alpha^{*})\geq 1$ per world (the problem is non-trivial—abnormality is necessary).
Gap constraint.
$\mathrm{cost}(\alpha^{*})\leq\mathrm{OptCost}+1$ per world, ensuring the planted gold is close to optimal.
Exception-rate cap.
Per-world abnormal fraction $\leq 0.20$:
$\frac{\mathrm{OptCost}(W)}{|D_{W}|}\;\leq\;0.20.$
This keeps “abnormality” a minority phenomenon as intended by default-exception logic.
Holdout worlds.
For each instance k=5k{=}5 holdout worlds are sampled from the same distribution and parameters, but without adversarial competitor elimination, to test generalisation under distributional shift.
D.3 Competitor Pool and CEGIS-like Elimination
Instances are hardened by adversarial world addition. The procedure ensures that no simple shortcut formula can achieve the same cost as the planted gold.
  1. 1.
    Initialise pool. Collect up to $C_{\max}=30$ competitor formulas:
    • Hand-curated Tier-1 shortcuts (e.g. $P(x)$, $\lnot Q(x)$, $\exists y\,R(x,y)$);
    • Mined Tier-2 shortcuts observed from model outputs;
    • Up to 10 syntactic mutants of the gold formula.
    Only competitors whose predicates are a subset of the allowed set are retained.
  2. 2.
    Test on current worlds. Evaluate every surviving competitor on the current training-world set. A competitor is beaten if it is invalid in $\geq 1$ world or its cost exceeds the gold cost by at least a margin $m$ (typically m=2m{=}2).
  3. 3.
    Add adversarial world. If survivors remain and the world budget is not exhausted, sample a new world that invalidates $\geq 1$ surviving competitor. Go to step 2.
  4. 4.
    Terminate. Accept the instance when all competitors are beaten. Reject if the world budget (typically 12–15) is exhausted with survivors remaining.
Algorithm D.1 gives the pseudocode.
function GENERATE_INSTANCE(theory, $\alpha*$, W_budget, m):
worlds = [sample_world(theory)]
pool = build_pool(theory, $\alpha*$, C_max)
while |worlds| < W_budget:
survivors = [c in pool :
is_valid(c, worlds) and
cost(c, worlds) < cost($\alpha*$, worlds) + m]
if survivors == []:
return ACCEPT(worlds)
w = sample_adversarial(theory, survivors)
if w is None: break // cannot separate
worlds.append(w)
return REJECT("survivors remain")
Figure D.1: CEGIS-like adversarial world addition.
D.4 Gold Refinement and Cheater Filtering
After competitor elimination, a second filter guards against cheater formulas—trivial patterns that happen to match the gold cost but have no semantic depth.
Cheater pool.
The cheater pool comprises two tiers:
Tier 1 (${\sim}25$ formulas):
Constants ($\top$, $\bot$), atomic unary/binary predicates and their negations, self-loops ($R(x,x)$, $S(x,x)$), simple quantified existence ($\exists y\,R(x,y)$), and unary conjunctions/disjunctions.
Tier 2 (${\sim}15$ formulas):
Common model shortcuts mined from preliminary runs, e.g. $\exists y(R(x,y)\land P(y))$, $P(x)\land\exists y(R(x,y))$.
Rejection criterion.
An instance is rejected if the best cheater achieves cost $\leq\mathrm{cost}(\alpha^{*})-1$ on the current training worlds. This ensures that the planted gold rule is not trivially beatable by a simple heuristic.
Gold refinement (optional).
When enabled, the pipeline evaluates ${\sim}20$ alternative gold candidates (sampled from the theory’s template library plus mutants of the seed) and selects the one with the best validity-cost-cheater-margin trade-off.
Appendix E Completion Semantics and Solver Encoding
All validity checks and cost computations are implemented via reduction to SMT using the Z3 solver. This section formalises the encoding.
E.1 World Representation
A world $W$ over domain $D_{W}=\{a_{0},\ldots,a_{n-1}\}$ is specified by:
  • $\mathrm{True}_{P}(W)$: set of atoms known true for each predicate $P$;
  • $\mathrm{Unk}_{P}(W)$: set of atoms whose truth value is unknown;
  • all remaining atoms are known false (closed-world assumption on the “rest”).
A completion $\mathcal{C}$ assigns each unknown atom a truth value. The completed world $W^{\mathcal{C}}$ has:
$\mathrm{Facts}(W^{\mathcal{C}})\;=\;\mathrm{True}_{P}(W)\;\cup\;\{\,a\in\mathrm{Unk}_{P}(W):\mathcal{C}(a)=\top\,\}.$
Z3 encoding.
Each unknown atom $u\in\mathrm{Unk}_{P}(W)$ is represented by a free Z3 boolean variable. Known-true atoms are replaced by True, known-false atoms by False, and unknown atoms by their variable. The abnormality predicate $\mathsf{Ab}(a)$ is replaced inline by the Z3 grounding of $\alpha(a)$, making the biconditional $\mathsf{Ab}(a)\leftrightarrow\alpha(a)$ implicit.
E.2 Validity Queries
Let $\Theta$ denote the theory axioms and $\Phi_{W}(\alpha,\mathcal{C})$ abbreviate the conjunction of grounded axioms after substituting $\mathsf{Ab}\mapsto\alpha$ and applying completion $\mathcal{C}$.
ABD-Full.
No unknown atoms exist. Validity reduces to direct evaluation:
$\mathrm{Valid}_{\mathrm{full}}(\alpha,W)\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;W\models\Theta[\mathsf{Ab}\mapsto\alpha].$
ABD-Partial (existential completion).
$\mathrm{Valid}_{\exists}(\alpha,W)\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\exists\,\mathcal{C}.\;W^{\mathcal{C}}\models\Theta[\mathsf{Ab}\mapsto\alpha].$
Encoding. Assert all grounded axioms with unknown atoms as free variables. The query is Sat; a satisfying assignment witnesses a valid completion.
ABD-Skeptical (universal completion).
$\mathrm{Valid}_{\forall}(\alpha,W)\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\forall\,\mathcal{C}.\;W^{\mathcal{C}}\models\Theta[\mathsf{Ab}\mapsto\alpha].$
Encoding. Assert the negation of the grounded axiom conjunction with unknown atoms as free variables. The query checks for a counterexample completion:
  • Unsat $\Rightarrow$ no counterexample exists $\Rightarrow$ $\alpha$ is valid.
  • Sat $\Rightarrow$ counterexample found $\Rightarrow$ $\alpha$ is invalid.
E.3 Cost Computation
The abnormality cost in world $W$ under completion $\mathcal{C}$ is
$\mathrm{cost}(\alpha,W^{\mathcal{C}})\;=\;\bigl|\{\,a\in D_{W}:W^{\mathcal{C}}\models\alpha(a)\,\}\bigr|.$
ABD-Full.
Direct evaluation; no optimisation needed.
ABD-Partial (best-case cost).
$\mathrm{cost}_{\exists}(\alpha,W)\;=\;\min_{\mathcal{C}:\;\mathrm{Valid}_{\exists}(\alpha,W^{\mathcal{C}})}\;\mathrm{cost}(\alpha,W^{\mathcal{C}}).$
Encoding. Use Z3’s Optimize solver: assert the grounded axioms (with unknown atoms as free variables) and minimise $\sum_{a\in D_{W}}\mathbf{1}[\alpha(a)]$.
ABD-Skeptical (worst-case cost).
$\mathrm{cost}_{\forall}(\alpha,W)\;=\;\max_{\mathcal{C}:\;\mathrm{Valid}_{\forall}(\alpha,W^{\mathcal{C}})}\;\mathrm{cost}(\alpha,W^{\mathcal{C}}).$
Encoding. Use Z3’s Optimize solver: assert the grounded axioms and maximise $\sum_{a\in D_{W}}\mathbf{1}[\alpha(a)]$. This finds the least favourable valid completion.
E.4 Baselines: Free-$\mathsf{Ab}$ Lower Bound
The conservative baseline $\mathrm{OptCost}$ relaxes the single-formula constraint:
ABD-Full.
$\mathrm{OptCost}(W)\;=\;\min_{\mathsf{Ab}:D_{W}\to\{0,1\}}\;\bigl|\{a:\mathsf{Ab}(a)=1\}\bigr|\quad\text{s.t.}\quad W\models\Theta.$
Here $\mathsf{Ab}$ is a free function (one independent boolean per domain element), not constrained to be definable by a single formula. The solution is computed via Z3 Optimize.
ABD-Partial.
$\mathrm{OptCost}_{\exists}(W)\;=\;\min_{\mathsf{Ab},\;\mathcal{C}}\;\bigl|\{a:\mathsf{Ab}(a)=1\}\bigr|\quad\text{s.t.}\quad W^{\mathcal{C}}\models\Theta.$
Both $\mathsf{Ab}$ and the completion $\mathcal{C}$ are jointly optimised.
ABD-Skeptical.
$\mathrm{OptCost}_{\forall}(W)\;=\;\min_{\mathsf{Ab}}\;\max_{\mathcal{C}}\;\bigl|\{a:\mathsf{Ab}(a)=1\}\bigr|\quad\text{s.t.}\quad\forall\,\mathcal{C}.\;W^{\mathcal{C}}\models\Theta.$
Conservativeness.
Because the single-formula definability constraint ($\mathsf{Ab}(a)\leftrightarrow\alpha(a)$ for a single $\alpha$) is dropped, $\mathrm{OptCost}$ is a lower bound on the achievable cost of any formula. The gap $\mathrm{Gap}=\mathrm{cost}(\alpha)-\mathrm{OptCost}$ therefore over-estimates actual sub-optimality but provides a consistent, formula-independent baseline across all models and instances.
Appendix F Holdout Construction and Distribution Match
F.1 Holdout Generation Algorithm
Each benchmark instance includes k=5k{=}5 holdout worlds used exclusively for post-hoc generalisation evaluation; no model ever sees holdout worlds during inference. The holdout sampler reuses the identical Z3-based world-generation procedure described in Appendix D.1:
  1. 1.
    Extract parameters from the training worlds of each instance: domain-size range, predicate-density intervals (Table D.1), unknown-atom rates (Table D.2), cost bounds, and gap constraints.
  2. 2.
    Sample a world using the same Z3 density-constrained generator. For ABD-Partial and ABD-Skeptical, unknown atoms are masked using the same per-theory rates.
  3. 3.
    Filter. Accept the world only if it (i) is not equivalent to any training world (identical domain size and predicate extensions), (ii) satisfies the validity semantics of the scenario, and (iii) has per-world cost and gap within the range observed across training worlds.
  4. 4.
    Repeat until $k$ distinct holdout worlds are accepted (up to 150 attempts per world, 30 s per instance).
This filtering ensures that the planted gold formula $\alpha^{*}$ remains valid on every holdout world and that per-world cost and gap fall within the training regime, so holdout evaluation measures generalisation under distributional match rather than distribution shift.
Deterministic seeding.
Each holdout world is seeded via $\texttt{SHA-256}(\mathit{dataset\_path}\mathbin{\|}\mathit{instance\_id}\mathbin{\|}\mathit{holdout\_idx}\mathbin{\|}\mathit{global\_seed})\bmod 2^{31}$, ensuring full reproducibility.
F.2 What Is Not Done on Holdout
The following training-time hardening procedures (Appendix D.3D.4) are deliberately omitted from holdout generation:
  • No competitor elimination. Training worlds are iteratively added to defeat shortcut formulas (§D.3); holdout worlds are drawn independently with no adversarial loop.
  • No cheater filtering. Training rejects instances where a trivial pattern matches the gold cost (§D.4); holdout applies no such filter.
  • No gold refinement. The planted gold formula is fixed at benchmark-creation time and not re-optimised for holdout.
These omissions are intentional: holdout worlds provide a pure i.i.d. distributional test, measuring whether a model’s formula captures the true pattern rather than artifacts of adversarial training-set construction.
F.3 Train vs. Holdout Distribution Match
Table F.1 compares per-world summary statistics across train and holdout sets. Per-world Gold Cost and OptCost are computed as $\mathrm{sum}/k$ where $k$ is the number of worlds. The close agreement confirms that holdout worlds are drawn from the same generating distribution.
Table F.1: Per-world summary statistics for train and holdout sets. Values are means across all instances in each scenario. Gold Gap = Gold Cost - OptCost, per world.
ABD-Full ABD-Partial ABD-Skeptical
Metric Train Hold Train Hold Train Hold
Worlds/inst 9.0 5.0 8.3 5.0 6.0 5.0
PW Gold Cost 1.78 1.88 1.68 1.75 2.05 2.10
PW OptCost 1.27 1.33 1.20 1.32 1.36 1.54
PW Gold Gap 0.51 0.55 0.48 0.43 0.69 0.56
Appendix G Additional Results
G.1 Full Per-Theory Results
Tables G.1G.3 report per-theory Train-Gap (T) and Holdout-Gap (H) for every model, with the number of survivors (instances valid on both train and holdout) in parentheses. Cells marked “—” indicate zero train validity. Gaps are normalised per world.
Table G.1: Per-theory results: ABD-Full. Each cell shows T-Gap / H-Gap (survivors).
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 1.3/3.2 (4) 1.1/3.2 (3) 1.3/3.7 (3) 1.5/3.6 (4) 1.1/1.9 (3) 1.4/3.2 (5) 1.5/3.3 (5) 1.4/3.1 (4)
T2 1.4/3.3 (56) 0.9/1.6 (35) 1.1/3.2 (37) 1.4/3.2 (54) 1.2/2.8 (38) 1.4/3.0 (49) 1.6/3.6 (48) 1.4/3.5 (33) 2.0/3.7 (6)
T3 1.4/2.8 (37) 1.1/2.5 (29) 0.9/2.6 (19) 1.3/2.6 (39) 1.4/2.6 (21) 1.6/2.9 (39) 1.6/2.9 (36) 1.6/3.0 (24) 1.4/3.0 (4)
T4 1.3/2.5 (46) 0.9/1.4 (27) 0.9/2.5 (28) 1.4/2.2 (44) 1.4/2.2 (25) 1.4/2.4 (49) 1.5/2.7 (45) 1.7/2.6 (24) 5.3/— (0)
T5 0.8/1.8 (27) 0.7/1.8 (22) 0.6/2.2 (18) 0.9/2.0 (27) 0.9/2.0 (25) 0.9/2.1 (24) 1.0/2.1 (26) 1.3/2.4 (24) 2.9/4.3 (6) 5.8/7.2 (1)
Table G.2: Per-theory results: ABD-Partial. Format as Table G.1.
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 1.0/2.5 (5) 1.3/3.2 (4) 0.9/3.3 (2) 1.0/3.0 (3) 2.3/2.4 (3) 1.4/3.0 (5) 1.4/3.0 (5) 1.3/2.9 (4)
T2 1.2/2.7 (38) 1.5/2.9 (56) 1.4/2.8 (52) 1.5/2.9 (56) 1.3/2.4 (28) 1.6/3.0 (67) 1.6/2.9 (63) 1.7/3.0 (54) 1.8/3.0 (2) 2.9/3.2 (9)
T3 1.2/2.1 (36) 1.5/2.7 (42) 1.3/2.4 (34) 1.3/2.3 (38) 1.2/2.0 (22) 1.6/2.6 (49) 1.5/2.5 (43) 1.7/2.6 (34) 1.7/2.6 (24)
T4 0.9/1.7 (22) 1.6/3.0 (43) 1.4/2.7 (48) 1.5/2.3 (37) 1.3/2.4 (21) 2.0/3.1 (57) 1.7/2.9 (63) 1.9/3.3 (41) 5.0/5.6 (4)
T5 0.6/1.4 (21) 1.4/2.4 (27) 0.9/1.9 (27) 1.0/1.9 (31) 0.9/2.1 (21) 1.4/2.4 (28) 1.2/2.1 (33) 1.3/2.4 (24) 5.8/— (0)
[…]
Opus-4.6 Grok4 GPT-5.2 Gemini-3.1 Gemini-3 Grok4.1f DSR Kimi-K2t Hermes4 GPT-4o
T1 1.0/1.8 (10) 1.2/2.4 (4) 1.4/2.3 (14) 1.2/2.3 (10) 1.0/2.0 (10) 1.8/2.0 (13) 1.8/2.5 (14) 2.7/2.3 (18) 4.0/4.0 (1) 5.5/9.2 (1)
T2 1.1/2.6 (14) 1.4/2.8 (11) 1.8/3.2 (14) 1.5/2.9 (16) 1.2/2.1 (14) 1.8/2.4 (16) 2.0/3.4 (16) 3.4/4.4 (18) 1.9/5.8 (1) 2.6/4.9 (7)
T3 1.4/2.4 (13) 1.3/2.4 (10) 1.7/3.1 (12) 1.4/2.2 (10) 1.2/1.8 (11) 1.8/2.7 (14) 1.7/2.7 (14) 3.2/4.3 (16) 2.0/4.4 (1) 2.8/4.3 (5)
T4 0.9/1.7 (13) 1.2/2.6 (11) 2.1/2.7 (17) 1.5/2.0 (15) 1.7/2.4 (14) 1.5/1.9 (8) 2.3/2.9 (24) 2.9/3.5 (19)
T5 1.1/1.0 (13) 1.6/1.3 (12) 2.4/1.3 (22) 1.7/1.2 (15) 1.8/1.3 (18) 2.0/1.6 (15) 2.5/1.5 (21) 2.5/1.5 (19) 4.9/4.2 (2) 6.9/— (0)
T6 0.2/0.1 (3) 0.2/0.1 (2) 0.2/0.7 (4) 0.2/0.1 (2) 0.2/0.3 (4) 0.5/1.3 (7) 0.3/0.8 (5) 0.2/0.0 (2) 2.9/3.2 (12)
T7 0.1/0.4 (13) 0.3/0.5 (12) 0.4/0.7 (16) 0.4/0.7 (16) 1.2/0.6 (12) 0.7/0.9 (19) 0.7/1.0 (19) 1.2/1.4 (21) 1.0/1.8 (1) 1.7/2.1 (15)
Key observations.
(i) Hermes4 and GPT-4o achieve near-zero train validity for most theories in ABD-Full and ABD-Partial, reflecting their difficulty with multi-world abduction. (ii) In ABD-Skeptical, holdout gaps can be lower than train gaps (T5), because worst-case cost under universal completion varies across world draws. (iii) Opus-4.6 achieves the highest survivor counts and near-100% train validity across all scenarios, with competitive gap levels; in ABD-Skeptical it achieves the tightest gap distribution (median 0.20, only 9.4% above gap 3).
G.2 Gap Distribution Summaries
Table G.4 reports the distribution of normalised train-set gap (cost above OptCost, per world) for each model. Only train-valid predictions are included.
Table G.4: Normalised train-gap distribution by scenario. N = predictions with all-valid on train. >>3, >>5 = tail percentages.
Model N Mean Med P90 Max >>3 >>5
ABD-Full
Opus-4.6 195 1.26 1.25 2.00 2.7 0.0% 0.0%
Grok4 171 0.92 0.90 1.62 2.4 0.0% 0.0%
GPT-5.2 166 0.93 0.95 2.00 2.4 0.0% 0.0%
Gemini-3.1 194 1.27 1.20 1.97 7.9 1.0% 1.0%
Gemini-3 151 1.24 1.12 2.10 8.2 1.3% 0.7%
Grok4.1f 183 1.35 1.30 2.17 3.0 0.0% 0.0%
DSR 183 1.45 1.40 2.20 3.2 0.5% 0.0%
Kimi-K2t 111 1.48 1.50 2.20 2.9 0.0% 0.0%
Hermes4 6 2.85 2.25 5.80 5.8 16.7% 16.7%
GPT-4o 17 3.05 1.88 5.90 6.5 41.2% 35.3%
ABD-Partial
Opus-4.6 238 1.02 1.00 1.67 2.9 0.0% 0.0%
Grok4 220 1.54 1.56 2.33 5.4 1.8% 0.5%
GPT-5.2 226 1.30 1.22 2.11 3.0 0.0% 0.0%
Gemini-3.1 236 1.41 1.22 2.13 7.0 2.5% 0.8%
Gemini-3 160 1.22 1.11 1.89 7.8 2.5% 0.6%
Grok4.1f 231 1.70 1.62 2.67 5.9 3.9% 0.4%
DSR 234 1.54 1.56 2.33 3.5 0.9% 0.0%
Kimi-K2t 176 1.69 1.62 2.62 4.0 3.4% 0.0%
Hermes4 2 1.78 1.78 2.33 2.3 0.0% 0.0%
GPT-4o 55 3.19 2.33 5.75 6.8 41.8% 23.6%
ABD-Skeptical
Opus-4.6 160 0.86 0.20 2.83 6.7 9.4% 1.9%
Grok4 144 1.08 0.33 3.17 6.7 10.4% 2.1%
GPT-5.2 162 1.49 0.63 4.00 8.0 20.4% 4.3%
Gemini-3.1 158 1.18 0.37 3.58 7.8 12.0% 3.8%
Gemini-3 150 1.25 0.43 3.83 6.7 15.3% 4.7%
Grok4.1f 128 1.46 0.71 3.83 8.0 17.2% 5.5%
DSR 151 1.73 1.33 4.40 6.8 21.9% 4.6%
Kimi-K2t 130 2.46 2.18 4.71 10.0 28.5% 6.2%
Hermes4 6 3.11 2.50 6.80 6.8 33.3% 16.7%
GPT-4o 47 2.87 2.20 5.80 7.4 27.7% 14.9%
Discussion.
In ABD-Full, the top six models show tightly concentrated gap distributions (P90 $\leq 2.3$, tail >3>3 at most 1.3%). ABD-Partial shows slightly heavier tails due to cost computation under existential completion. ABD-Skeptical exhibits the widest gap distributions: weaker models have over 20% of predictions with gap >3>3 (worst-case cost under universal completion inflates gaps), while Opus-4.6 achieves the tightest distribution (median 0.20, P90 2.83, only 9.4% above gap 3).
G.3 Beats-Gold Breakdown
A prediction beats gold when its total train cost is strictly below the planted gold formula’s cost on the same training worlds. This is possible because the gold formula is a fixed template that may not be optimal for every world draw.
By theory.
Table G.5 aggregates beats-gold rates across all models.
Table G.5: Beats-gold rate by theory (all models pooled). $k/N$ = predictions beating gold out of train-valid predictions.
Theory ABD-Full ABD-Partial ABD-Skeptical
T1 0/30  (0.0%) 1/28  (3.6%) 9/143  (6.3%)
T2 31/318  (9.7%) 4/422  (0.9%) 13/123 (10.6%)
T3 22/225  (9.8%) 4/310  (1.3%) 26/126 (20.6%)
T4 23/309  (7.4%) 5/401  (1.2%) 32/140 (22.9%)
T5 32/175 (18.3%) 8/185  (4.3%) 23/141 (16.3%)
T6 48/83  (57.8%)
T7 5/138  (3.6%)
T6 in ABD-Skeptical stands out at 57.8%: the planted gold for TH3 ($P(x)\land\lnot\mathsf{Ab}(x)\to\exists y\,R(x,y)$) yields a generous gold cost under worst-case completion, and models often find tighter formulas. ABD-Partial has universally low beats-gold rates ($\leq 4.3\%$), reflecting good gold-formula design for existential completion.
By gold AST size.
Table G.6 groups beats-gold rate by the AST size of the planted gold formula.
Table G.6: Beats-gold rate by gold formula AST size (all models pooled).
Gold AST ABD-Full ABD-Partial ABD-Skeptical
$[0,5)$ 0/6  (0.0%)
$[5,10)$ 24/161 (14.9%)
$[10,15)$ 19/132 (14.4%) 9/370  (2.4%) 52/397 (13.1%)
$[15,20)$ 51/397 (12.8%) 6/438  (1.4%) 33/123 (26.8%)
$[20,30)$ 10/406  (2.5%) 5/420  (1.2%) 44/183 (24.0%)
$[30,+)$ 28/122 (23.0%) 1/113  (0.9%) 3/24  (12.5%)
In ABD-Full, the highest beats-gold rate occurs for the most complex golds (AST $\geq 30$, 23.0%), where the planted formula contains redundancy that models can trim. In ABD-Skeptical, mid-range golds (AST 15–30) are beaten most often (24–27%), suggesting these formulas are complex enough to leave room for simpler alternatives under worst-case semantics.
Top beats-gold examples.
The largest single-instance margin is +14+14 in ABD-Skeptical (instance abd_skeptical_v2_TH3_00, T6): five models independently discover a formula with train cost 7 vs. gold cost 21. In ABD-Full, the largest margin is +7+7 (GPT-5.2 on ABD_FULL_TH7_021, T2: predicted cost 12 vs. gold cost 19). ABD-Partial margins are smaller (max +4+4), consistent with the tight gold design under existential completion.
Appendix H Qualitative Examples
H.1 Worked Examples per Regime
For each observation regime we select one instance that has both a robust prediction (valid on train and holdout) and a brittle prediction (valid on train, fails on holdout), illustrating the generalisation challenge. Formulas are shown in S-expression notation.
ABD-Full: instance ABD_FULL_TH7_021 (T2).
Theory T2: $\exists y(R(x,y)\land P(y))\land\lnot\mathsf{Ab}(x)\to\exists z(S(x,z)\land Q(z))$. 10 training worlds, 5 holdout worlds. 5 models produce robust formulas; 2 are brittle.
Robust (Grok4.1f, AST = 37):
(exists y (exists z (and (not (= y z))
  (R x y) (R x z) (or (P y) (P z))
  (forall w (or (not (R x w)) (or (= w y) (= w z)))))))
Train cost = 13 (gap 0.10/world); holdout cost = 9, all valid.
Interpretation: $x$ has exactly two $R$-successors, at least one of which satisfies $P$. The cardinality constraint generalises because the domain structure is stable.
Brittle (Gemini-3, AST = 22):
(exists y (and (R x y) (P y) (forall z (or (not (and (R x z) (P z))) (= y z)))))
Train cost = 30; holdout: 3/5 worlds invalid (40% valid).
Failure: insists on a unique $P$-witness ($\forall z$-clause forces $z=y$). Three holdout worlds happen to have two or more $R$-related $P$-elements, violating uniqueness.
ABD-Partial: instance ABD_PARTIAL_TH7_014 (T2).
Same theory as above, but under partial observation. 8 training worlds, 5 holdout worlds. 1 robust, 6 brittle predictions.
Robust (GPT-4o, AST = 8):
(exists y (and (R x y) (P y)))
Train cost = 22 (gap 1.50/world); holdout cost = 23, all valid.
Interpretation: the simplest sufficient condition—$x$ has an $R$-related $P$-element. Generalises precisely because it avoids over-specification.
Brittle (Gemini-3, AST = 12):
(and (exists y (and (R x y) (P y))) (not (P x)))
Train cost = 17; holdout: 0/5 worlds valid (catastrophic failure).
Failure: the conjunct $\lnot P(x)$ happens to hold for all abnormal elements in the 8 training worlds, but is not a consequence of the theory. All 5 holdout worlds contain $P$-positive abnormals, breaking the formula.
ABD-Skeptical: instance abd_skeptical_v2_TH2_11 (T1).
Theory T1: $\exists y(R(x,y)\land P(y))\land\lnot\mathsf{Ab}(x)\to Q(x)$. 5 training worlds, 5 holdout worlds. 3 robust, 3 brittle predictions.
Robust (Opus-4.6, AST = 18):
(and (P x) (exists y (R x y)) (forall z (or (not (R x z)) (P z))))
Train cost = 5 (gap 0.00/world); holdout cost = 5, all valid.
Interpretation: $x$ is abnormal if $P(x)$ holds, $x$ has some $R$-successor, and all $R$-successors of $x$ satisfy $P$. The universal quantifier over $R$-successors makes this robust under worst-case completion of unknown $R$-atoms.
Brittle (GPT-5.2, AST = 19):
(exists y (and (R x y) (P y) (forall z (or (not (R x z)) (= z y)))))
Train cost = 5; holdout: catastrophic failure.
Failure: the uniqueness constraint ($\forall z$-clause forces $z=y$) insists that $x$ has exactly one $R$-related $P$-element. Holdout worlds readily contain elements with multiple $P$-witnesses, invalidating the formula under skeptical completion.
H.2 Common Failure Patterns
Table H.1 gives the per-model failure breakdown aggregated across all tasks.
Table H.1: Failure Mode Breakdown. Aggregated across all abduction tasks. Parse = formula extraction failed; Invalid = parsed but not valid on all train worlds; Valid = train-valid predictions. Best values bold.
Model Valid% Invalid% Parse% Missing%
Opus-4.6 98.8% 0.5% 0.0% 0.7%
Grok4 89.3% 10.7% 0.0% 0.0%
GPT-5.2 92.3% 7.7% 0.0% 0.0%
Gemini-3.1 98.0% 2.0% 0.0% 0.0%
Gemini-3 76.8% 23.2% 0.0% 0.0%
Grok4.1f 94.8% 5.2% 0.0% 0.0%
DSR 94.7% 5.3% 0.0% 0.0%
Kimi-K2t 69.5% 30.5% 0.0% 0.0%
Hermes4 2.3% 97.7% 0.0% 0.0%
GPT-4o 19.8% 80.2% 0.0% 0.0%
Table H.2 categorises all model predictions by failure mode. A prediction is classified into the first matching category:
  1. 1.
    Parse error: formula does not parse as a valid S-expression.
  2. 2.
    All invalid (train): valid on zero training worlds.
  3. 3.
    Partial invalid (train): valid on some but not all training worlds.
  4. 4.
    Brittle: valid on all training worlds, invalid on $\geq 1$ holdout world. Sub-category catastrophic: $50\%$ holdout worlds valid.
  5. 5.
[…]
  6. 6.
[…]
Table H.2: Failure-mode counts across all 5,971 model predictions.
Failure Mode ABD-Full ABD-Partial ABD-Skeptical
Parse error 0 4 0
All invalid (train) 332 406 315
Partial invalid (train) 241 241 41
Brittle (holdout) 254 452 465
of which catastrophic 73 99 265
Parsimony inflation 306 189 72
Observations.
Parse errors are negligible (4 total, all from ABD-Partial). The dominant failure mode in ABD-Partial is all invalid on train (406), reflecting the difficulty of satisfying existential completion across 8+ worlds. Brittle failures increase monotonically with scenario difficulty: ABD-Skeptical has the most brittle predictions (465) and the highest catastrophic rate (265/465 = 57% of brittle predictions fail on $\geq$50% of holdout worlds, vs. 29% for ABD-Full). Parsimony inflation moves in the opposite direction (306 in Full, 72 in Skeptical): models in Skeptical tend to fail validity outright rather than produce valid-but-expensive formulas.
Brittle formula patterns.
Table H.3 groups brittle predictions by syntactic pattern.
Table H.3: Brittle formula syntactic patterns. “Exists-only” = existential quantifiers without universals.
Pattern ABD-Full ABD-Partial ABD-Skeptical
Exists-only 126 215 225
Mixed quantifiers 79 32 47
Propositional (and) 0 9 19
Forall-only 0 3 9
Other 1 10 2
Exists-only formulas dominate across all regimes: they assert the existence of witnesses with specific properties that hold incidentally on training worlds (e.g. a unique $P$-witness, a self-loop, a specific neighbourhood count). These formulas tend to be brittle because existential witnesses are sensitive to the particular world draw. The mixed quantifiers pattern is second most common in ABD-Full, where models attempt precise cardinality constraints (“exactly two $R$-successors”) that break on holdout worlds with different domain structure.
Appendix I Software and Data
To avoid compromising anonymity during an ongoing review process, we do not publicly release the benchmark artifacts at this time. After the review process, we will release the frozen v1 datasets and results, the generation/evaluation/analysis pipeline, and cached model outputs sufficient to reproduce all reported tables and figures.